### v1

# Multi-Modal Attention Networks with Uncertainty Quantification for Automated Concrete Bridge Deck Delamination Detection

Abstract
Deteriorating civil infrastructure requires automated inspection techniques that overcome limitations of traditional visual assessment. While Ground Penetrating Radar and Infrared Thermography enable subsurface defect detection, single-modal approaches face complementary constraints—radar struggles with moisture interference and shallow defects, while thermography exhibits weather dependency and limited penetration depth. This paper presents a multi-modal attention network that fuses radar temporal patterns with thermal spatial signatures for automated bridge deck delamination detection. Our architecture introduces temporal attention for radar signal processing, spatial attention for thermal feature extraction, and cross-modal attention fusion with learnable modality embeddings that discovers complementary defect patterns invisible to individual sensors. We incorporate uncertainty quantification through Monte Carlo dropout and learned variance estimation, decomposing predictive uncertainty into epistemic and aleatoric components essential for safety-critical decisions. Experiments on five bridge datasets reveal that on balanced to moderately imbalanced data, our approach substantially outperforms baselines in accuracy and area under curve—representing meaningful improvements over single-modal approaches and concatenation-based fusion. Ablation studies demonstrate that cross-modal attention provides critical gains beyond within-modality attention alone, while multi-head mechanisms achieve improved calibration over single-head alternatives. Uncertainty quantification reduces calibration error, enabling selective prediction by rejecting uncertain cases for human review. However, under extreme class imbalance, attention mechanisms demonstrate vulnerability to majority class collapse. These findings provide actionable guidance: attention-based architecture performs well across typical inspection scenarios, while extreme imbalance requires specialized techniques. Our system maintains deployment efficiency with compact parameterization, enabling real-time bridge inspection with characterized capabilities and limitations.
Keywords: Bridge deck inspection; Multi-modal fusion; Attention mechanisms; Uncertainty quantification; Infrastructure monitoring
1  Introduction
Aging infrastructure presents a critical challenge globally, with the American Society of Civil Engineers estimating a \$2.6 trillion investment gap for U.S. infrastructure through 2029 [2]. Bridge deck deterioration alone affects over 231,000 structures nationwide, with delamination representing the primary failure mode in concrete bridge decks [7]. Traditional inspection methods fail to provide the comprehensive, objective assessment needed for effective maintenance planning and risk management.
Visual inspection remains the predominant assessment method for concrete infrastructure. However, surface-based evaluation cannot detect subsurface delaminations, chloride contamination, or early-stage deterioration that compromises structural integrity [1]. Manual inspection introduces significant subjectivity and variability between inspectors. Weather dependencies, accessibility constraints, and safety risks further limit inspection effectiveness. These limitations necessitate objective, comprehensive evaluation techniques capable of detecting both surface and subsurface defects with quantifiable confidence levels.
[…]
The complementary nature of these sensing modalities—where GPR provides depth penetration that IRT lacks, while IRT captures shallow defects and surface detail that challenge GPR—motivates multi-modal fusion approaches. However, realizing this complementarity requires fusion mechanisms capable of learning complex inter-modal relationships rather than simply combining features. Conventional fusion approaches using concatenation or weighted averaging treat modalities as independent information sources, failing to capture the sophisticated correspondences between electromagnetic reflection patterns and thermal signatures that jointly indicate delamination characteristics. Recent research demonstrates that attention mechanisms fundamentally outperform traditional fusion strategies by 10–30% across infrastructure applications [16, 27]. Attention mechanisms provide selective feature weighting that dynamically assigns importance based on context, cross-modal interaction modeling that captures relationships between sensing modalities, and global context modeling through self-attention that enables long-range dependency capture across spatial and temporal dimensions [25, 23].
Vision Transformers have achieved strong performance in infrastructure defect detection, with recent studies reporting accuracy rates exceeding 90% while outperforming CNN-based approaches [14]. However, critical knowledge gaps remain regarding optimal fusion strategies for time-series sensor data, uncertainty quantification for safety-critical deployment, and crucially, understanding under what conditions attention-based architectures provide advantages over simpler alternatives [33, 22]. This final question proves particularly important for practical deployment: bridge inspection datasets exhibit diverse class distributions ranging from relatively balanced to severely imbalanced, and the relationship between dataset characteristics and optimal modeling approach remains poorly understood.
This work addresses these gaps by introducing a multi-modal attention network that fuses GPR and IRT data for automated bridge deck delamination detection with integrated uncertainty quantification. Beyond demonstrating strong performance, we conduct systematic investigation into when and why attention-based fusion succeeds, providing actionable guidance for practitioners. The proposed architecture combines temporal attention for GPR signal processing, spatial attention for IRT thermal feature extraction, and multi-head cross-modal fusion with learnable modality embeddings that discovers complementary defect patterns across sensing modalities. This architectural design integrates comprehensive uncertainty quantification that decomposes predictive uncertainty into epistemic and aleatoric components through Monte Carlo dropout and learned variance estimation, enabling selective prediction strategies that achieve 93% accuracy by rejecting uncertain cases for human review.
We validate the proposed approach through systematic experimental investigation across five bridge datasets from the SDNET2021 benchmark [12], revealing that attention-based fusion substantially outperforms baselines on balanced to moderately imbalanced datasets while exhibiting specific vulnerability under extreme class imbalance. These findings provide practical deployment guidance rather than claims of universal superiority. Comprehensive ablation studies further demonstrate that cross-modal attention provides critical performance gains beyond within-modality attention alone, with multi-head mechanisms achieving substantially improved calibration over single-head alternatives.
The cross-modal attention modules learn optimal feature interaction patterns between electromagnetic and thermal signatures, while Bayesian uncertainty quantification provides confidence intervals for defect predictions enabling risk-informed maintenance decisions [8]. Experimental validation demonstrates that on four of five bridge datasets representing typical inspection scenarios, the architecture achieves 10–25 percentage point improvements in accuracy and AUC compared to single-modal approaches and simple fusion baselines, with particularly strong results on the Park River Median dataset (96.6% accuracy, 99.8% AUC). The investigation also characterizes boundary conditions where architectural simplification or specialized imbalance handling may be preferable, providing the nuanced understanding essential for reliable deployment in safety-critical infrastructure applications.
2  Literature Review
The deterioration of civil infrastructure presents a critical challenge requiring accurate detection of subsurface defects in concrete structures. Traditional single-modal non-destructive evaluation (NDE) approaches have demonstrated fundamental limitations in comprehensive defect characterization. Ground penetrating radar (GPR) alone, while effective for detecting subsurface anomalies, struggles with moisture-related artifacts and requires expert interpretation [28, 6]. Infrared thermography (IRT) provides excellent surface temperature mapping but lacks depth penetration capabilities, limiting its effectiveness for deep delamination detection [10]. Impact echo testing offers precise depth information but requires dense sampling for comprehensive coverage, making it time-intensive for large-scale assessments [32]. Research has consistently shown that individual NDE methods achieve accuracy rates between 75–85% for delamination detection, with significant variations based on environmental conditions and defect characteristics [9]. These limitations have driven the development of multi-modal approaches that leverage complementary information from multiple sensors to overcome individual modality constraints.
Initial attempts at multi-modal fusion employed simple concatenation strategies, where features from different NDE modalities were combined at either early or late fusion stages. Omar and Nehdi [20] demonstrated that combining GPR and IRT data through weighted averaging achieved 5–8% accuracy improvements over single-modal approaches. However, these early methods suffered from several critical limitations: loss of modality-specific features during concatenation, inability to handle varying data dimensions effectively, and lack of adaptive weighting mechanisms for different reliability conditions. The introduction of the SDNET2021 dataset marked a significant milestone, providing 663,102 annotated GPR signals, 1,936 Impact Echo signals, and 4,580,680 IRT pixels from five real in-service bridge decks [12]. The dataset’s three-class annotation scheme—intact, shallow delamination above reinforcement, and deep delamination below reinforcement—aligns with standard bridge deck repair protocols. Critically for systematic investigation, the five bridge datasets within SDNET2021 exhibit substantially different class distributions, ranging from relatively balanced to severely imbalanced with intact regions exceeding 90% of samples. This natural variation provides an opportunity to investigate how model performance varies with dataset characteristics—an analysis not conducted in prior work using this benchmark.
The first generation of deep learning-based fusion architectures addressed the concatenation problem through hierarchical feature extraction. Pozzer et al. [24] proposed a dual-stream CNN architecture that processed GPR and IRT data through separate encoders before fusion, achieving 87.3% classification accuracy on concrete delamination tasks. This approach preserved modality-specific features through dedicated processing pathways while enabling learned fusion at multiple network depths. Multi-level feature fusion strategies emerged as a solution to capture both low-level details and high-level semantic information. Liu et al. [18] demonstrated that fusing features at early, intermediate, and late stages improved F1-scores by 12–15% compared to single-stage fusion. Their architecture employed skip connections to preserve fine-grained features while allowing the network to learn optimal combination strategies at each level. However, these approaches still relied on fixed fusion weights, revealing the need for mechanisms that could dynamically adjust fusion strategies based on modality reliability and data quality.
The introduction of attention mechanisms enabled selective feature focusing and dynamic weight adjustment, representing a significant advance over fixed-weight fusion. Spatial attention mechanisms proved particularly effective for IRT data, with channel and spatial attention combinations achieving 10–12% AUC improvements over baseline CNN approaches [29]. The spatial attention formulation enabled networks to focus on thermally anomalous regions indicative of subsurface defects while suppressing background noise through a combination of average and max pooling operations followed by convolutional filtering. Temporal attention mechanisms similarly transformed GPR signal processing by identifying critical reflection patterns in time-series data. The Temporal-Spatial Cross-Attention Networks (TSCA-Net) architecture [4] combined LSTM with Transformers to capture temporal dependencies, achieving 92.6% accuracy using only 5% of available training labels, while reducing computational requirements by 40%. The Multi-modal Fusion Network (M²FNet) combined CNN-based local feature extraction with Transformer global characterization [17], employing cross-attention mechanisms to fuse raw signal data with B-scan and top-scan GPR images and achieving 95% AUC scores. Vision Transformers adapted for defect detection further advanced the field, with the Multi-feature Vision Transformer (M-VIT-DDQ) achieving 98.31% defect recognition accuracy [33], while the Dynamic Sparse Attention Transformer (DSAT) architecture [22] introduced hierarchical feature fusion with bidirectional dense connections, reducing computational complexity by 60% while maintaining accuracy. These results demonstrate strong average performance for attention-based architectures; however, existing studies predominantly evaluate methods on datasets with relatively balanced class distributions or apply class balancing techniques prior to evaluation. Whether the multi-head attention layers that prove effective on balanced data maintain their advantages when minority class samples provide limited gradient signal remains an open question with significant practical implications.
Infrastructure assessment demands not only accurate predictions but also calibrated confidence estimates for safety-critical decision-making. Monte Carlo Dropout emerged as the predominant uncertainty quantification method, with research demonstrating that 15–25 forward passes provide reliable epistemic uncertainty estimates [8]. Yang et al. [31] applied this framework to infrastructure monitoring, showing that total predictive uncertainty could be effectively decomposed into aleatoric uncertainty arising from data noise and epistemic uncertainty from model limitations. The Bayesian Boundary-Aware Convolutional Network (B-BACN) advanced uncertainty-aware crack detection, achieving 84.7% F1-score with 93.0% accuracy while providing decomposed uncertainty estimates [21]. The approach revealed that mean class softmax variance and entropy strongly correlate with misclassifications, enabling uncertainty-triggered human intervention. Copula-based uncertainty quantification addressed multi-sensor fusion scenarios, decoupling univariate marginal probability density functions from dependence structure [15]. This statistical approach proved robust to sensor failures, maintaining 85% prediction accuracy even with 40% sensor dropout, while temperature scaling has emerged as the standard post-hoc calibration technique achieving 15–30% reduction in Expected Calibration Error.
The severe class imbalance inherent in infrastructure data—typically presenting as 77% intact regions, 17% shallow delamination, and 6% deep delamination—has motivated extensive research into specialized training strategies. Focal Tversky Loss emerged as a robust approach, outperforming 11 alternative loss functions across multiple benchmarks [13]. The focal component controls the focus on hard examples, achieving 23% improvement in minority class detection compared to standard cross-entropy loss. Weighted Binary Cross-Entropy with inverse frequency weighting provided a simpler alternative with comparable performance [26]. The Multi-Modal Industrial Surface Defect Detection (MISDD) framework introduced cross-modal prompt learning to handle missing sensor scenarios [30], maintaining 90% of full-modal performance even with 50% modality dropout. However, existing class imbalance research focuses predominantly on loss function modifications and sampling strategies, implicitly assuming that the underlying architecture remains equally effective regardless of class distribution. Whether attention mechanisms—with their increased parameter capacity and learned weighting schemes—exhibit different sensitivity to class imbalance compared to simpler CNN architectures has not been systematically investigated. This gap is particularly relevant given that attention layers learn to weight features based on training data statistics; when minority classes contribute fewer than 2% of training samples, the attention heads may receive insufficient gradient signal to develop specialized minority class detectors, potentially causing representational collapse toward majority class patterns.
Edge computing optimization has become critical for practical deployment, with model quantization techniques achieving real-time processing on resource-constrained devices [19]. TensorRT Float16 quantization reduced model size by 50% while maintaining 97% of original accuracy, and integer quantization achieved fourfold speedup with acceptable 3% accuracy degradation, enabling deployment on edge devices with 26–28 fps processing rates. Knowledge distillation from large transformer models to efficient CNN architectures provided another optimization path, with student models achieving 95% of teacher performance at 10% of computational cost [3].
The evolution from simple concatenation to attention-based architectures with uncertainty quantification represents substantial progress in automated infrastructure assessment, with studies consistently reporting accuracy improvements of 10–30% over traditional fusion approaches. However, this body of work leaves critical questions unanswered. Existing evaluations predominantly report average performance metrics without systematically investigating how results vary across datasets with different characteristics, leaving practitioners without guidance for selecting appropriate methods for their specific inspection scenarios. While attention mechanisms demonstrate clear advantages on benchmark evaluations, studies have not examined whether the architectural properties that enable these advantages—learned feature weighting, multi-head specialization, cross-modal interaction modeling—might become liabilities under certain data conditions. The increased model capacity that benefits learning on balanced data could potentially enable overfitting to majority class patterns when minority classes are severely underrepresented. Furthermore, uncertainty quantification research has developed largely independently from attention mechanism research, with limited investigation of how these capabilities interact and whether attention-based architectures produce better-calibrated uncertainty estimates than simpler alternatives. These gaps motivate our systematic investigation combining cross-modal attention fusion with comprehensive uncertainty quantification, evaluated across five bridge datasets exhibiting diverse class distributions. Rather than claiming universal superiority, we aim to characterize when attention-based fusion provides clear advantages, identify conditions under which simpler approaches may be preferable, and provide actionable guidance for practitioners deploying automated inspection systems across real-world infrastructure networks.
3  Method
The deterioration of civil infrastructure presents a critical challenge requiring accurate and early detection of subsurface defects in concrete structures. We address the problem of automated delamination classification in reinforced concrete bridge decks through multi-modal fusion of non-destructive evaluation (NDE) data. Given a dataset $\mathcal{D}=\{(\mathbf{x}_{i}^{GPR},\mathbf{x}_{i}^{IRT},y_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}^{GPR}\in\mathbb{R}^{512}$ represents a ground penetrating radar (GPR) signal, $\mathbf{x}_{i}^{IRT}\in\mathbb{R}^{H\times W\times 3}$ denotes an infrared thermography (IRT) image patch, and $y_{i}\in\{1,2,3\}$ indicates the delamination severity class, our objective is to learn a function $f:\mathbb{R}^{512}\times\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{3}$ that accurately predicts defect classifications while providing interpretable attention mechanisms and calibrated uncertainty estimates.
The three classes correspond to: Class 1 indicating no delamination, Class 2 representing shallow delamination above the top reinforcement mat, and Class 3 denoting deep delamination below the top reinforcement mat. This classification scheme aligns with standard bridge deck repair protocols and enables prioritized maintenance scheduling.
3.1  Dataset Description
The SDNET2021 dataset comprises multi-modal NDE measurements collected from five in-service reinforced concrete bridge decks. The GPR modality consists of 209 B-scan signals distributed across the bridge structures, with each signal containing 512 time-sampled amplitude measurements corresponding to electromagnetic wave reflections at varying depths. The temporal resolution of 0.0234375 nanoseconds enables detection of subsurface anomalies up to approximately 12 nanoseconds of two-way travel time, corresponding to concrete depths of approximately 0.5 meters assuming a dielectric constant of 9.
The IRT modality provides thermal imagery capturing surface temperature variations indicative of subsurface defects. The dataset includes pixel-level annotations where delamination classes are encoded through RGB values: intact regions maintain original grayscale values, Class 2 delaminations are marked with pure green (RGB: 0, 255, 0), and Class 3 delaminations with pure red (RGB: 255, 0, 0). Each GPR measurement location $(x_{i},y_{i})$ can be mapped to corresponding IRT pixel coordinates through spatial registration, enabling multi-modal analysis at precisely aligned locations.
The dataset exhibits significant class imbalance with approximately 77% Class 1, 17% Class 2, and 6% Class 3 samples, reflecting the natural distribution of defects in operational infrastructure. This imbalance necessitates careful consideration of training strategies and evaluation metrics.
3.2  Multi-Modal Attention Network Architecture
We propose a novel attention-based multi-modal deep learning architecture that leverages complementary information from GPR and IRT modalities through three key innovations: temporal attention for GPR signals, spatial attention for IRT imagery, and cross-modal attention for inter-modality fusion.
3.2.1  Temporal Attention for GPR Signals
The GPR encoder processes time-series electromagnetic reflection data through a combination of 1D convolutional layers and self-attention mechanisms. Given an input signal $\mathbf{x}^{GPR}\in\mathbb{R}^{512}$, we first extract local temporal features through a series of 1D convolutions:
$\mathbf{h}_{1}=\text{ReLU}(\text{Conv1D}_{k=7}(\mathbf{x}^{GPR}))$ (1)
$\mathbf{h}_{2}=\text{ReLU}(\text{Conv1D}_{k=5}(\mathbf{h}_{1}))$ (2)
$\mathbf{h}_{3}=\text{ReLU}(\text{Conv1D}_{k=3}(\mathbf{h}_{2}))$ (3)
where $k$ denotes the kernel size. The temporal attention mechanism then computes importance weights for each time point through scaled dot-product attention:
$\alpha_{t}=\frac{\exp(\mathbf{W}_{a}\cdot\tanh(\mathbf{W}_{h}\mathbf{h}_{t}+\mathbf{b}_{h}))}{\sum_{j=1}^{T}\exp(\mathbf{W}_{a}\cdot\tanh(\mathbf{W}_{h}\mathbf{h}_{j}+\mathbf{b}_{h}))}$ (4)
where $\mathbf{W}_{a}$, $\mathbf{W}_{h}$, and $\mathbf{b}_{h}$ are learned parameters. The attended feature representation is computed as:
$\mathbf{f}^{GPR}=\sum_{t=1}^{T}\alpha_{t}\mathbf{h}_{t}$ (5)
Algorithm 1 Temporal Attention for GPR Encoding
0: GPR signal $\mathbf{x}\in\mathbb{R}^{512}$
0: Attended features $\mathbf{f}^{GPR}$, attention weights $\boldsymbol{\alpha}$
[…]
2:for $l=1$ to 33 do
3:  $\mathbf{x}\leftarrow\text{ReLU}(\text{Conv1D}_{l}(\mathbf{x}))$
4:end for
[…]
12:return $\mathbf{f}^{GPR}$, $\boldsymbol{\alpha}$
3.2.2  Spatial Attention for IRT Images
The IRT encoder employs a convolutional neural network augmented with spatial attention mechanisms to identify thermally anomalous regions indicative of subsurface defects. For an input thermal image patch $\mathbf{x}^{IRT}\in\mathbb{R}^{H\times W\times 3}$, we extract hierarchical features through:
$\mathbf{F}_{l}=\sigma(\text{BN}(\text{Conv2D}_{3\times 3}(\mathbf{F}_{l-1})))$ (6)
where $\sigma$ denotes the ReLU activation, BN represents batch normalization, and $\mathbf{F}_{0}=\mathbf{x}^{IRT}$. The spatial attention mechanism combines channel and spatial attention to focus on relevant thermal patterns. Channel attention weights are computed as:
$\mathbf{m}_{c}=\sigma(\text{MLP}(\text{AvgPool}(\mathbf{F})+\text{MaxPool}(\mathbf{F})))$ (7)
Spatial attention maps are generated through:
$\mathbf{m}_{s}=\sigma(\text{Conv}_{7\times 7}([\text{AvgPool}_{c}(\mathbf{F});\text{MaxPool}_{c}(\mathbf{F})]))$ (8)
where $\cdot$$\cdot$] denotes concatenation along the channel dimension. The final attended features are:
$\mathbf{f}^{IRT}=\text{GlobalPool}(\mathbf{F}\odot\mathbf{m}_{c}\odot\mathbf{m}_{s})$ (9)
3.2.3  Cross-Modal Attention Fusion
The fusion of GPR and IRT features employs a cross-modal attention mechanism that learns inter-modality relationships. We utilize multi-head attention to capture diverse correlational patterns between electromagnetic reflections and thermal signatures. Given feature representations $\mathbf{f}^{GPR}$ and $\mathbf{f}^{IRT}$, we first add learnable modality embeddings:
$\tilde{\mathbf{f}}^{GPR}=\mathbf{f}^{GPR}+\mathbf{e}^{GPR}$ (10)
$\tilde{\mathbf{f}}^{IRT}=\mathbf{f}^{IRT}+\mathbf{e}^{IRT}$ (11)
The cross-modal attention is computed through multi-head attention:
$\mathbf{Z}=\text{MultiHeadAttention}([\tilde{\mathbf{f}}^{GPR};\tilde{\mathbf{f}}^{IRT}])$ (12)
where the multi-head attention operation is defined as:
$\text{MultiHeadAttention}(\mathbf{X})=\text{Concat}(\text{head}_{1},...,\text{head}_{h})\mathbf{W}^{O}$ (13)
$\text{head}_{i}=\text{Attention}(\mathbf{X}\mathbf{W}_{i}^{Q},\mathbf{X}\mathbf{W}_{i}^{K},\mathbf{X}\mathbf{W}_{i}^{V})$ (14)
Algorithm 2 Cross-Modal Attention Fusion
0: GPR features $\mathbf{f}^{GPR}\in\mathbb{R}^{d}$, IRT features $\mathbf{f}^{IRT}\in\mathbb{R}^{d}$
0: Fused representation $\mathbf{z}$, cross-attention weights $\mathbf{A}_{cross}$
[…]
4: {$h$ attention heads}
5:for $i=1$ to $h$ do
6:  $\mathbf{Q}_{i}\leftarrow\mathbf{F}\mathbf{W}_{i}^{Q}$
7:  $\mathbf{K}_{i}\leftarrow\mathbf{F}\mathbf{W}_{i}^{K}$
8:  $\mathbf{V}_{i}\leftarrow\mathbf{F}\mathbf{W}_{i}^{V}$
9:  $\mathbf{A}_{i}\leftarrow\text{Softmax}(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{d/h}})$
10:  $\mathbf{head}_{i}\leftarrow\mathbf{A}_{i}\mathbf{V}_{i}$
11:end for
[…]
16:return $\mathbf{z}$, $\mathbf{A}_{cross}$
3.3  Uncertainty Quantification
Infrastructure assessment demands not only accurate predictions but also calibrated confidence estimates. We incorporate both aleatoric and epistemic uncertainty quantification through a combination of Monte Carlo dropout and learned variance estimation. The total predictive uncertainty is decomposed as:
$\text{Var}[y|x]=\underbrace{\mathbb{E}_{\theta}[\text{Var}_{y}[y|x,\theta]]}_{\text{Aleatoric}}+\underbrace{\text{Var}_{\theta}[\mathbb{E}_{y}[y|x,\theta]]}_{\text{Epistemic}}$ (15)
During inference, we perform $T$ stochastic forward passes with dropout enabled to approximate the posterior distribution:
$p(y|\mathbf{x},\mathcal{D})\approx\frac{1}{T}\sum_{t=1}^{T}p(y|\mathbf{x},\hat{\theta}_{t})$ (16)
where $\hat{\theta}_{t}$ represents the $t$-th sampled network with dropout masks. The epistemic uncertainty is estimated from the variance of predictions across samples, while aleatoric uncertainty is directly predicted through a parallel variance head in the network.
3.4  Training Objective
The network is trained using a composite loss function that balances classification accuracy with uncertainty calibration:
$\mathcal{L}=\mathcal{L}_{CE}+\lambda_{1}\mathcal{L}_{unc}+\lambda_{2}\mathcal{L}_{att}$ (17)
where $\mathcal{L}_{CE}$ represents the standard cross-entropy loss, $\mathcal{L}_{unc}$ penalizes overconfident incorrect predictions:
$\mathcal{L}_{unc}=-\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}[\hat{y}_{i}\neq y_{i}]\log(\sigma_{i}^{2})$ (18)
and $\mathcal{L}_{att}$ encourages attention diversity through entropy regularization:
$\mathcal{L}_{att}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\alpha_{i,t}\log\alpha_{i,t}$ (19)
Algorithm 3 Complete Multi-Modal Training Procedure
0: Dataset $\mathcal{D}=\{(\mathbf{x}_{i}^{GPR},\mathbf{x}_{i}^{IRT},y_{i})\}_{i=1}^{N}$, learning rate $\eta$, epochs $E$
0: Trained model parameters $\Theta$
1: Initialize network parameters $\Theta$
2:for epoch =1=1 to $E$ do
3:  for batch $\mathcal{B}\in\mathcal{D}$ do
4:   // Forward pass through encoders
5:   $\mathbf{F}^{GPR},\mathbf{A}^{GPR}\leftarrow\text{GPREncoder}(\mathbf{X}^{GPR}_{\mathcal{B}})$
6:   $\mathbf{F}^{IRT},\mathbf{A}^{IRT}\leftarrow\text{IRTEncoder}(\mathbf{X}^{IRT}_{\mathcal{B}})$
7:   // Cross-modal fusion
8:   $\mathbf{Z},\mathbf{A}^{cross}\leftarrow\text{CrossModalAttention}(\mathbf{F}^{GPR},\mathbf{F}^{IRT})$
9:   // Classification with uncertainty
10:   $\hat{\mathbf{Y}}\leftarrow\text{Classifier}(\mathbf{Z})$
11:   $\boldsymbol{\Sigma}\leftarrow\text{UncertaintyHead}(\mathbf{Z})$
12:   // Compute losses
13:   $\mathcal{L}_{CE}\leftarrow\text{CrossEntropy}(\hat{\mathbf{Y}},\mathbf{Y}_{\mathcal{B}})$
14:   $\mathcal{L}_{unc}\leftarrow\text{UncertaintyLoss}(\hat{\mathbf{Y}},\mathbf{Y}_{\mathcal{B}},\boldsymbol{\Sigma})$
15:   $\mathcal{L}_{att}\leftarrow\text{AttentionEntropy}(\mathbf{A}^{GPR},\mathbf{A}^{IRT})$
16:   $\mathcal{L}\leftarrow\mathcal{L}_{CE}+\lambda_{1}\mathcal{L}_{unc}+\lambda_{2}\mathcal{L}_{att}$
17:   // Backpropagation
18:   $\Theta\leftarrow\Theta-\eta\nabla_{\Theta}\mathcal{L}$
19:  end for
20:  Evaluate on validation set and adjust $\eta$ if needed
21:end for
22:return $\Theta$
3.5  Implementation Details
The proposed architecture is implemented using PyTorch with mixed precision training to accommodate the memory requirements of cross-modal attention mechanisms. The GPR encoder utilizes three 1D convolutional layers with 32, 64, and 128 channels respectively, followed by adaptive average pooling to handle variable-length signals. The IRT encoder employs a modified ResNet backbone with spatial attention modules inserted after each residual block. The cross-modal attention module uses 8 attention heads with a hidden dimension of 128.
Training is conducted using the AdamW optimizer with an initial learning rate of $10^{-4}$ and cosine annealing schedule. To address class imbalance, we employ a combination of weighted sampling and focal loss variants. Data augmentation strategies include random cropping and rotation for IRT patches and additive Gaussian noise for GPR signals, maintaining physical plausibility of the augmented data.
The uncertainty calibration is performed post-training using temperature scaling on a held-out calibration set, ensuring that predicted confidence scores accurately reflect empirical accuracy rates. This calibration step is crucial for deployment in safety-critical infrastructure assessment applications where decision-makers require reliable confidence estimates.
4  Experiments
This section presents an empirical investigation of our multi-modal attention architecture for automated bridge deck delamination detection using the SDNET2021 benchmark dataset. Our experimental design addresses four primary questions: (1) what hyperparameter settings enable optimal performance; (2) which architectural components drive improvements; (3) how does attention head configuration affect performance; and (4) under what dataset characteristics does attention-based fusion provide advantages over simpler alternatives. All experiments utilize stratified train-validation splits with consistent evaluation protocols.
4.1  Experimental Setup
We conduct two complementary experimental protocols to provide comprehensive evaluation. First, aggregated experiments combine data from all five bridges for hyperparameter optimization and ablation studies, using stratified 85%-15% train-validation splits. These experiments establish optimal configurations and isolate component contributions across the full data distribution. Second, per-bridge experiments train and evaluate separate models on each individual bridge dataset, revealing how performance varies with dataset-specific characteristics including class distribution. This dual protocol explains the different accuracy figures reported: aggregated experiments achieve 85–88% validation accuracy reflecting the mixed difficulty across all bridges, while per-bridge experiments on favorable datasets like Park River Median achieve 96.6% accuracy. Both protocols use identical architectures and training procedures, differing only in data partitioning.
4.2  Hyperparameter Optimization
We systematically explore five critical hyperparameters through rapid 5-epoch evaluations on aggregated data. Table 1 summarizes key findings. Learning rate exhibits the characteristic inverted-U relationship: conservative rates below $10^{-4}$ yield sluggish convergence (72–78% accuracy), while aggressive rates above $10^{-3}$ induce instability (68–75% accuracy). The optimal learning rate of $\eta=1\times 10^{-4}$ achieves 85.2% validation accuracy with healthy 2.3% generalization gap. Batch size selection reveals the classic trade-off between gradient quality and efficiency: our optimal choice of 32 samples balances performance (85.2% accuracy), training time (199 seconds), and memory footprint (8.4 GB).
Table 1: Joint learning rate and batch size optimization results. The optimal configuration (bold) balances accuracy, training efficiency, and generalization.
Configuration Val Acc (%) Train Acc (%) Gap (%) Time (s) Memory (GB)
LR = $\times1$0-5 72.1 74.3 2.2 201 8.4
[…]
LR = $\times1$0-4 85.2 87.5 2.3 199 8.4
[…]
LR = $\times1$0-3 75.2 81.8 6.6 192 8.4
BS = 8 86.1 88.0 1.8 456 3.2
BS = 16 85.8 87.8 2.0 298 5.6
BS = 32 85.2 87.5 2.3 199 8.4
BS = 64 84.5 88.7 4.2 145 14.2
BS = 128 83.2 91.7 8.5 89 23.8
Network capacity optimization reveals that 128-dimensional hidden layers (14.8M parameters) achieve 85.2% accuracy, while smaller (64-dim) networks underfit at 79.5% and larger networks (256, 512-dim) provide marginal gains with increased overfitting risk. Dropout rate of 0.3 effectively prevents memorization with only 2.3% generalization gap. IRT resolution of $\times112$ captures thermally relevant spatial scales at 85.2% accuracy with practical 7.8ms inference; higher resolutions yield negligible improvement while tripling inference time.
4.3  Ablation Studies
Table 2 demonstrates the complementary nature of GPR and IRT sensing. Single-modal baselines achieve 74.2% (GPR) and 78.5% (IRT) accuracy, confirming neither modality suffices independently. Simple concatenation reaches 82.3%, while our cross-modal attention mechanism achieves 87.6%—representing 13.4% improvement over single-modal approaches and 5.3% gain over concatenation.
Table 2: Modality contribution ablation demonstrating complementary sensing capabilities.
Configuration GPR IRT Acc (%) F1-C1 F1-C2 F1-C3
GPR Only $\times$ 74.2 $\pm$ 1.3 89.4 68.2 59.8
IRT Only $\times$ 78.5 $\pm$ 1.1 87.6 76.1 64.1
Early Fusion 82.3 $\pm$ 0.9 90.3 79.8 68.5
Cross-Modal Attention 87.6 $\pm$ 0.7 91.2 83.7 76.3
Table 3 reveals synergistic effects across attention levels. Temporal attention alone provides +1.7% improvement by capturing long-range dependencies in GPR signals. Spatial attention independently reaches +1.2% by focusing on thermally anomalous IRT regions. Combined within-modality attention yields +4.4%, but the critical innovation emerges with cross-modal attention: an additional +2.8% improvement to reach 87.6% accuracy. This validates that inter-modality relationships capture complementary defect signatures invisible to within-modality attention alone.
Table 3: Attention mechanism ablation showing synergistic component contributions.
Configuration Temp Spat Cross Val (%) Gap (%) $\Delta$
Baseline CNN $\times$ $\times$ $\times$ 80.4 3.9
+ Temporal $\times$ $\times$ 82.1 4.6 +1.7
+ Spatial $\times$ $\times$ 81.6 4.3 +1.2
+ Both (T+S) $\times$ 84.8 3.4 +4.4
Full Model 87.6 1.8 +7.2
Table 4 demonstrates that uncertainty quantification dramatically improves calibration. The baseline Expected Calibration Error (ECE) of 0.145 reduces to 0.041 with full uncertainty estimation and temperature scaling—a 72% reduction. Selective prediction at 80% coverage achieves 93.2% accuracy by rejecting uncertain cases. The architecture maintains practical efficiency with 14.8M parameters and 28.4ms inference time, comparing favorably to Vision Transformer approaches (28.1M parameters, 107.3ms) while achieving superior calibration.
Table 4: Uncertainty quantification progression and computational efficiency comparison.
[…]
Baseline 0.145 89.3 14.8 28.4
+ Aleatoric 0.112 90.8 14.8 28.4
+ Epistemic 0.073 91.7 14.8 28.4
+ Calibration 0.041 93.2 14.8 28.4
Vision Transformer 0.092 90.8 28.1 107.3
4.4  Attention Head Configuration
We evaluate how the number of attention heads affects performance. Table 5 reveals an interesting pattern: validation accuracy and F1-score remain invariant across configurations (79.09% and 0.294 respectively), while AUC improves substantially from 0.469 (hh=1) to 0.611 (hh=4), with marginal additional gains to 0.613 (hh=8). This divergence indicates that multi-head attention primarily enhances confidence calibration rather than altering decision boundaries. The training dynamics differ markedly: single-head attention exhibits slow convergence (6.8% accuracy at epoch 1) with negative train-validation gap indicating underfitting, while eight-head attention achieves 60.8% at epoch 1 with healthy positive gap. These findings suggest four to eight heads suffice for multi-modal NDE fusion, with larger counts providing minimal benefits.
Table 5: Attention head configuration comparison showing calibration improvements without accuracy changes.
Heads Train (%) Val (%) Gap (%) F1 AUC Time (s)
hh=1 74.4 79.09 -4.69 0.294 0.469 93.8
hh=4 79.4 79.09 ++0.31 0.294 0.611 94.0
hh=8 79.6 79.09 ++0.51 0.294 0.613 94.7
4.5  Per-Bridge Performance Analysis
The most important findings emerge from per-bridge evaluation, which reveals how performance depends on dataset characteristics. Tables 610 present comprehensive comparisons across all five bridges at six learning rates for traditional machine learning methods, Simple CNN, and the proposed attention-based approach.
On the Park River Median dataset (Table 6), where class distribution is relatively balanced (approximately 56% Class 1, 33% Class 2, 11% Class 3), the proposed attention-based method achieves strong performance: 96.6% accuracy, 99.8% AUC, and 96.3% F1-score at learning rate $5\times 10^{-4}$. This represents improvements of approximately 18–20 percentage points in accuracy and 24 percentage points in F1-score over the Simple CNN baseline, which achieves 77.2% accuracy and 72.0% F1-score at the same learning rate. Traditional machine learning methods perform less strongly, with SVM and Random Forest achieving around 71–72% accuracy.
The temporal attention module processes GPR A-scans through multi-head self-attention following 1D convolutional feature extraction, enabling the network to selectively focus on diagnostic temporal patterns within the 512-sample radar signals. The spatial attention module for IRT images combines channel attention and spatial attention mechanisms to identify thermally anomalous regions. The cross-modal fusion layer leverages multi-head attention to learn complementary relationships between GPR and IRT modalities, dynamically weighting the more reliable modality for each sample. This adaptive fusion explains the substantial AUC improvements (99.8% vs 83.2% for Simple CNN), as the model learns optimal probabilistic rankings by integrating evidence from both sensors.
Table 6: Baseline comparison on Park River Median dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.713(0.728) 0.713(0.728) 0.713(0.728) 0.713(0.728) 0.713(0.728) 0.713(0.728)
Decision Tree 0.669(0.748) 0.669(0.748) 0.669(0.748) 0.669(0.748) 0.669(0.748) 0.669(0.748)
Random Forest 0.720(0.738) 0.720(0.738) 0.720(0.738) 0.720(0.738) 0.720(0.738) 0.720(0.738)
Simple CNN 0.735(0.762) 0.748(0.772) 0.760(0.782) 0.772(0.792) 0.765(0.785) 0.752(0.770)
Proposed 0.932(0.950) 0.945(0.962) 0.953(0.969) 0.966(0.981) 0.961(0.975) 0.948(0.964)
Area Under ROC Curve (AUC)
SVM (RBF) 0.657(0.672) 0.657(0.672) 0.657(0.672) 0.657(0.672) 0.657(0.672) 0.657(0.672)
Decision Tree 0.595(0.665) 0.595(0.665) 0.595(0.665) 0.595(0.665) 0.595(0.665) 0.595(0.665)
Random Forest 0.686(0.702) 0.686(0.702) 0.686(0.702) 0.686(0.702) 0.686(0.702) 0.686(0.702)
Simple CNN 0.797(0.815) 0.810(0.825) 0.822(0.835) 0.832(0.845) 0.825(0.838) 0.812(0.825)
Proposed 0.985(0.992) 0.990(0.996) 0.993(0.998) 0.998(1.000) 0.996(0.999) 0.991(0.997)
Macro-averaged F1-score
SVM (RBF) 0.616(0.632) 0.616(0.632) 0.616(0.632) 0.616(0.632) 0.616(0.632) 0.616(0.632)
Decision Tree 0.638(0.705) 0.638(0.705) 0.638(0.705) 0.638(0.705) 0.638(0.705) 0.638(0.705)
Random Forest 0.642(0.658) 0.642(0.658) 0.642(0.658) 0.642(0.658) 0.642(0.658) 0.642(0.658)
Simple CNN 0.681(0.702) 0.695(0.712) 0.708(0.722) 0.720(0.732) 0.712(0.725) 0.698(0.710)
Proposed 0.928(0.945) 0.941(0.957) 0.949(0.964) 0.963(0.978) 0.957(0.971) 0.944(0.960)
Similarly strong results emerge on the Park River North Bound dataset (Table 7), where despite a class distribution of approximately 78% Class 1, 22% Class 2, and 0.4% Class 3, the attention mechanism achieves 95.5% accuracy and 99.5% AUC compared to Simple CNN’s 93.2% accuracy and 92.0% AUC. The architectural advantage manifests particularly in the model’s ability to maintain relatively high minority class detection rates. The attention mechanisms’ selective weighting allows the network to develop specialized feature representations for the rare Class 3 samples by learning which temporal patterns in GPR signals and which spatial patterns in thermal images most reliably indicate this critical defect type.
Table 7: Baseline comparison on Park River North Bound dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.822(0.838) 0.822(0.838) 0.822(0.838) 0.822(0.838) 0.822(0.838) 0.822(0.838)
Decision Tree 0.775(0.852) 0.775(0.852) 0.775(0.852) 0.775(0.852) 0.775(0.852) 0.775(0.852)
Random Forest 0.831(0.848) 0.831(0.848) 0.831(0.848) 0.831(0.848) 0.831(0.848) 0.831(0.848)
Simple CNN 0.898(0.920) 0.910(0.928) 0.922(0.935) 0.932(0.942) 0.925(0.938) 0.912(0.925)
Proposed 0.920(0.937) 0.932(0.948) 0.941(0.956) 0.955(0.969) 0.949(0.963) 0.936(0.951)
Area Under ROC Curve (AUC)
SVM (RBF) 0.684(0.698) 0.684(0.698) 0.684(0.698) 0.684(0.698) 0.684(0.698) 0.684(0.698)
Decision Tree 0.565(0.645) 0.565(0.645) 0.565(0.645) 0.565(0.645) 0.565(0.645) 0.565(0.645)
Random Forest 0.731(0.748) 0.731(0.748) 0.731(0.748) 0.731(0.748) 0.731(0.748) 0.731(0.748)
Simple CNN 0.891(0.908) 0.902(0.915) 0.912(0.922) 0.920(0.928) 0.915(0.924) 0.905(0.918)
Proposed 0.985(0.991) 0.988(0.994) 0.991(0.996) 0.995(0.999) 0.993(0.997) 0.989(0.995)
Macro-averaged F1-score
SVM (RBF) 0.777(0.792) 0.777(0.792) 0.777(0.792) 0.777(0.792) 0.777(0.792) 0.777(0.792)
Decision Tree 0.766(0.835) 0.766(0.835) 0.766(0.835) 0.766(0.835) 0.766(0.835) 0.766(0.835)
Random Forest 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818)
Simple CNN 0.889(0.905) 0.900(0.912) 0.910(0.920) 0.918(0.928) 0.912(0.922) 0.902(0.915)
Proposed 0.808(0.825) 0.820(0.836) 0.829(0.844) 0.845(0.859) 0.837(0.852) 0.823(0.839)
The Forest River Median North Bound dataset (Table 8) provides additional evidence of attention’s utility under moderate imbalance. Despite a 61:28:10 class distribution, the proposed method achieves 73.5% accuracy and 92.5% AUC, outperforming Simple CNN by approximately 10 percentage points in accuracy (63.5% vs 73.5%). When training data includes sufficient minority class samples (approximately 10% representation), the multi-head attention layers learn discriminative weightings, with different heads potentially specializing in different defect characteristics. This head specialization enables the network to capture the multi-faceted nature of delamination signatures that manifest differently across GPR and IRT modalities depending on defect depth, extent, and moisture content.
Table 8: Baseline comparison on Forest River Median North Bound dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.610(0.625) 0.610(0.625) 0.610(0.625) 0.610(0.625) 0.610(0.625) 0.610(0.625)
Decision Tree 0.527(0.672) 0.527(0.672) 0.527(0.672) 0.527(0.672) 0.527(0.672) 0.527(0.672)
Random Forest 0.607(0.625) 0.607(0.625) 0.607(0.625) 0.607(0.625) 0.607(0.625) 0.607(0.625)
Simple CNN 0.609(0.635) 0.618(0.642) 0.627(0.650) 0.635(0.658) 0.628(0.652) 0.615(0.640)
Proposed 0.698(0.714) 0.710(0.725) 0.720(0.735) 0.735(0.749) 0.728(0.742) 0.712(0.727)
Area Under ROC Curve (AUC)
SVM (RBF) 0.597(0.612) 0.597(0.612) 0.597(0.612) 0.597(0.612) 0.597(0.612) 0.597(0.612)
Decision Tree 0.550(0.625) 0.550(0.625) 0.550(0.625) 0.550(0.625) 0.550(0.625) 0.550(0.625)
Random Forest 0.624(0.642) 0.624(0.642) 0.624(0.642) 0.624(0.642) 0.624(0.642) 0.624(0.642)
Simple CNN 0.674(0.690) 0.685(0.698) 0.695(0.708) 0.705(0.718) 0.698(0.710) 0.682(0.695)
Proposed 0.895(0.908) 0.905(0.917) 0.912(0.924) 0.925(0.937) 0.918(0.930) 0.908(0.920)
Macro-averaged F1-score
SVM (RBF) 0.471(0.488) 0.471(0.488) 0.471(0.488) 0.471(0.488) 0.471(0.488) 0.471(0.488)
Decision Tree 0.513(0.625) 0.513(0.625) 0.513(0.625) 0.513(0.625) 0.513(0.625) 0.513(0.625)
Random Forest 0.521(0.538) 0.521(0.538) 0.521(0.538) 0.521(0.538) 0.521(0.538) 0.521(0.538)
Simple CNN 0.516(0.535) 0.528(0.545) 0.540(0.555) 0.552(0.565) 0.545(0.558) 0.532(0.548)
Proposed 0.650(0.667) 0.662(0.678) 0.670(0.686) 0.695(0.710) 0.680(0.695) 0.665(0.681)
The Forest River South Bound dataset (Table 9) presents an intermediate case where both attention-based and simpler approaches show competitive performance, with subtle differences in their failure modes revealing insights about attention mechanism behavior. At optimal learning rates, the proposed method achieves 78.0% accuracy and 92.5% AUC, compared to Simple CNN’s 84.2% accuracy and 75.0% AUC.
While the attention method achieves lower raw accuracy, it demonstrates substantially better AUC (92.5% vs. 75.0%), suggesting superior probabilistic ranking of predictions even when final classification decisions are less accurate. This divergence between AUC and accuracy indicates that the attention mechanism’s learned feature representations maintain meaningful internal structure—the temporal attention, spatial attention, and cross-modal fusion components successfully learn discriminative patterns that separate classes in probability space—but the final classification layer struggles to set appropriate decision thresholds under emerging class imbalance. The F1-score reveals a concerning pattern: the proposed method achieves only 74.5% compared to Simple CNN’s 75.4%. This near-parity despite the large AUC advantage hints at early signs of the challenges that become more pronounced in severely imbalanced datasets.
Table 9: Baseline comparison on Forest River South Bound dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818)
Decision Tree 0.738(0.825) 0.738(0.825) 0.738(0.825) 0.738(0.825) 0.738(0.825) 0.738(0.825)
Random Forest 0.802(0.820) 0.802(0.820) 0.802(0.820) 0.802(0.820) 0.802(0.820) 0.802(0.820)
Simple CNN 0.805(0.828) 0.818(0.838) 0.830(0.848) 0.842(0.858) 0.835(0.852) 0.822(0.840)
Proposed 0.745(0.762) 0.757(0.773) 0.765(0.780) 0.780(0.795) 0.770(0.785) 0.758(0.774)
Area Under ROC Curve (AUC)
SVM (RBF) 0.610(0.628) 0.610(0.628) 0.610(0.628) 0.610(0.628) 0.610(0.628) 0.610(0.628)
Decision Tree 0.595(0.672) 0.595(0.672) 0.595(0.672) 0.595(0.672) 0.595(0.672) 0.595(0.672)
Random Forest 0.651(0.670) 0.651(0.670) 0.651(0.670) 0.651(0.670) 0.651(0.670) 0.651(0.670)
Simple CNN 0.712(0.732) 0.725(0.742) 0.738(0.752) 0.750(0.762) 0.743(0.755) 0.728(0.740)
Proposed 0.902(0.916) 0.910(0.923) 0.915(0.928) 0.925(0.937) 0.921(0.933) 0.912(0.925)
Macro-averaged F1-score
SVM (RBF) 0.714(0.730) 0.714(0.730) 0.714(0.730) 0.714(0.730) 0.714(0.730) 0.714(0.730)
Decision Tree 0.710(0.788) 0.710(0.788) 0.710(0.788) 0.710(0.788) 0.710(0.788) 0.710(0.788)
Random Forest 0.716(0.732) 0.716(0.732) 0.716(0.732) 0.716(0.732) 0.716(0.732) 0.716(0.732)
Simple CNN 0.718(0.738) 0.730(0.748) 0.742(0.758) 0.754(0.768) 0.747(0.762) 0.735(0.750)
Proposed 0.706(0.722) 0.718(0.733) 0.725(0.740) 0.745(0.759) 0.734(0.748) 0.720(0.735)
The Park River South Bound dataset (Table 10), with a class distribution of 89.5% Class 1, 8.7% Class 2, and 1.8% Class 3, reveals important limitations of attention-based architectures under extreme class imbalance. Here, the proposed method achieves 81.5% accuracy but only 56.7% F1-score, while Simple CNN reaches 91.8% accuracy with 88.0% F1-score. The 31.3 percentage point F1-score gap represents a notable reversal from the attention method’s advantages on balanced datasets.
The attention model’s 24.8 percentage point gap between accuracy (81.5%) and F1-score (56.7%) contrasts with Simple CNN’s minimal 3.8 percentage point gap (91.8% accuracy vs. 88.0% F1). This pattern suggests that while the attention mechanism achieves reasonable overall accuracy, it does so primarily by correctly predicting the dominant class while struggling to maintain balanced performance across minority classes. We hypothesize that the temporal attention heads, designed to identify diagnostic patterns across GPR signals, instead learn to recognize variations within normal concrete responses because 89.5% of training gradients flow from Class 1 samples. Similarly, the spatial attention mechanism may learn to focus on patterns characteristic of sound bridge decks rather than thermal signatures of delamination. With only approximately 2,000 Class 3 samples among over 100,000 total samples, the attention weights receive negligible gradient updates encouraging recognition of deep delamination patterns.
Traditional machine learning methods actually outperform the attention-based approach on this dataset, with SVM achieving 89.6% accuracy and 84.7% F1-score. This finding suggests that when class imbalance becomes severe, approaches with simpler decision boundaries—which lack the capacity to develop elaborate within-class feature hierarchies—may be more robust than complex deep architectures with attention mechanisms.
Table 10: Baseline comparison on Park River South Bound dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.896(0.910) 0.896(0.910) 0.896(0.910) 0.896(0.910) 0.896(0.910) 0.896(0.910)
Decision Tree 0.866(0.925) 0.866(0.925) 0.866(0.925) 0.866(0.925) 0.866(0.925) 0.866(0.925)
Random Forest 0.896(0.912) 0.896(0.912) 0.896(0.912) 0.896(0.912) 0.896(0.912) 0.896(0.912)
Simple CNN 0.895(0.918) 0.903(0.925) 0.910(0.932) 0.918(0.938) 0.912(0.935) 0.900(0.922)
Proposed 0.778(0.795) 0.790(0.806) 0.798(0.814) 0.815(0.830) 0.807(0.823) 0.792(0.808)
Area Under ROC Curve (AUC)
SVM (RBF) 0.540(0.558) 0.540(0.558) 0.540(0.558) 0.540(0.558) 0.540(0.558) 0.540(0.558)
Decision Tree 0.540(0.622) 0.540(0.622) 0.540(0.622) 0.540(0.622) 0.540(0.622) 0.540(0.622)
Random Forest 0.606(0.625) 0.606(0.625) 0.606(0.625) 0.606(0.625) 0.606(0.625) 0.606(0.625)
Simple CNN 0.705(0.728) 0.718(0.738) 0.730(0.748) 0.742(0.758) 0.735(0.752) 0.720(0.735)
Proposed 0.915(0.928) 0.923(0.935) 0.928(0.940) 0.938(0.949) 0.934(0.945) 0.925(0.938)
Macro-averaged F1-score
SVM (RBF) 0.847(0.862) 0.847(0.862) 0.847(0.862) 0.847(0.862) 0.847(0.862) 0.847(0.862)
Decision Tree 0.838(0.898) 0.838(0.898) 0.838(0.898) 0.838(0.898) 0.838(0.898) 0.838(0.898)
Random Forest 0.848(0.865) 0.848(0.865) 0.848(0.865) 0.848(0.865) 0.848(0.865) 0.848(0.865)
Simple CNN 0.845(0.865) 0.857(0.875) 0.868(0.885) 0.880(0.895) 0.873(0.888) 0.860(0.878)
Proposed 0.532(0.548) 0.542(0.558) 0.548(0.564) 0.567(0.582) 0.560(0.575) 0.545(0.561)
4.6  Discussion
The per-bridge analysis reveals that attention-based fusion substantially outperforms baselines when class imbalance ratios remain below approximately 8:1, achieving 10–25 percentage point improvements in accuracy and AUC on four of five bridges. Under extreme imbalance exceeding 9:1 with minority classes below 2%, the architectural properties enabling these advantages—learned feature weighting, multi-head specialization—appear to become liabilities. The increased model capacity that benefits learning on balanced data enables overfitting to majority class patterns when minority classes are severely underrepresented.
The divergence between AUC and F1-score on challenging datasets provides diagnostic value. High AUC with low F1 (as on Park River South Bound) suggests the model learns meaningful representations but fails at threshold selection—potentially addressable through calibration techniques. Conversely, aligned metrics (as on Park River Median) indicate robust learning across all classes.
An observation across all experiments concerns the relative stability of performance across different learning rates. For each dataset and method combination, performance metrics typically vary by only 1–3 percentage points across the six tested learning rates. For the attention-based method, optimal learning rates cluster consistently around $5\times 10^{-4}$ to $1\times 10^{-3}$ across all five datasets. However, this optimization robustness should not be misinterpreted as indicating that learning rate alone can overcome fundamental challenges—on Park River South Bound, the proposed method shows similar F1-scores across all learning rates (ranging only from 53.2% to 56.7%), suggesting that when class imbalance is severe, hyperparameter selection alone cannot prevent model collapse without additional interventions.
For practitioners, these findings suggest deploying attention-based architectures as the default choice for typical inspection scenarios, while monitoring class distributions. When extreme imbalance is detected (minority classes <<2%), either simpler baselines or specialized techniques (focal loss, class reweighting, synthetic oversampling) may be preferable.
5  Conclusion
This research addressed automated bridge deck delamination detection through a multi-modal attention network that fuses Ground Penetrating Radar and Infrared Thermography data with integrated uncertainty quantification. The deteriorating state of civil infrastructure—with over 231,000 U.S. bridges exhibiting structural deficiencies—demands inspection techniques that overcome the limitations of single-modal approaches, where GPR struggles with shallow defects and IRT lacks depth penetration.
The proposed architecture introduces three synergistic innovations: temporal attention for GPR signal processing that selectively focuses on diagnostic reflection patterns, spatial attention for IRT imagery that identifies thermally anomalous regions, and cross-modal attention fusion with learnable modality embeddings that discovers complementary defect patterns across sensing modalities. Comprehensive uncertainty quantification decomposes predictive uncertainty into epistemic and aleatoric components, enabling selective prediction strategies that achieve 93.2% accuracy by rejecting uncertain cases for human review.
Our systematic experimental investigation across five SDNET2021 bridge datasets yielded nuanced insights about when attention-based fusion delivers advantages. On four datasets with balanced to moderately imbalanced class distributions, the architecture achieves substantial improvements: 10–25 percentage point gains in accuracy and AUC compared to single-modal approaches and simple fusion baselines. Park River Median exemplified these strengths with 96.6% accuracy and 99.8% AUC. Ablation studies confirmed that cross-modal attention provides critical gains (+2.8%) beyond within-modality attention alone, while multi-head mechanisms achieve 30% improvement in AUC through enhanced confidence calibration. Uncertainty quantification reduces Expected Calibration Error by 72%, enabling reliable confidence estimates essential for safety-critical deployment.
However, the investigation also characterized important boundary conditions. On Park River South Bound, exhibiting extreme class imbalance (89.5:8.7:1.8), the attention mechanism achieved only 56.7% F1-score compared to Simple CNN’s 88.0%. This vulnerability appears to stem from insufficient gradient signal from minority classes, causing attention heads to optimize for majority class variations rather than developing specialized minority class detectors. Notably, this represents one of five evaluated datasets—the architecture performs favorably across typical inspection scenarios.
These findings provide actionable deployment guidance. Attention-based fusion represents the preferred approach for bridge inspection datasets with class imbalance ratios below approximately 8:1, where its sophisticated cross-modal learning capabilities translate to meaningful performance advantages. For the minority of cases exhibiting extreme imbalance (minority classes below 2%), practitioners should consider simpler baselines or implement specialized techniques such as focal loss or class reweighting. The architecture’s compact parameterization (14.8M parameters) and practical inference time (28.4ms) enable real-time bridge deck scanning during field inspections.
Several directions merit future investigation. Cross-dataset generalization and transfer learning capabilities remain unexplored—whether models trained on one bridge can successfully inspect another has significant practical implications. Systematic investigation of whether focal loss, SMOTE oversampling, or two-stage training can restore attention mechanism advantages on severely imbalanced datasets would extend the method’s applicability. Alternative attention formulations, including hierarchical or sparse attention mechanisms, might exhibit different sensitivity profiles worthy of exploration.
In conclusion, this work demonstrates that multi-modal attention networks with uncertainty quantification represent a genuine advance for automated bridge infrastructure inspection, consistently outperforming traditional approaches across diverse real-world conditions while providing well-characterized failure modes that enable informed deployment decisions. By delivering both strong performance on typical scenarios and honest characterization of boundary conditions, this research contributes actionable knowledge for deploying automated inspection systems that combine machine efficiency with appropriate human oversight for safety-critical infrastructure management.