### v3

# Multi-Sensor Attention Networks for Automated Subsurface Delamination Detection in Concrete Bridge Decks

Abstract
Subsurface delaminations in concrete bridge decks remain undetectable through conventional visual inspection, necessitating automated non-destructive evaluation methods. This work introduces a deep learning framework that integrates Ground Penetrating Radar (GPR) and Infrared Thermography (IRT) through hierarchical attention mechanisms. Our architecture employs temporal self-attention to process GPR electromagnetic signals, spatial attention to analyze thermal imagery, and cross-modal attention with learnable embeddings to model inter-sensor correspondences. We integrate Monte Carlo dropout-based uncertainty quantification, decomposing prediction confidence into model uncertainty and data-driven uncertainty components. Testing across five real-world bridge datasets from the SDNET2021 benchmark reveals that our approach delivers substantial performance gains over single-sensor and concatenation-based baselines when applied to balanced or moderately imbalanced data distributions. Comprehensive ablation analysis confirms that cross-modal attention mechanisms contribute meaningful improvements beyond unimodal attention alone. Critically, we identify and characterize specific failure modes: under extreme class imbalance, attention-based architectures demonstrate susceptibility to majority class bias, indicating scenarios where simpler architectural choices may prove more robust. Our findings equip practitioners with empirically-grounded criteria for selecting appropriate fusion strategies based on dataset characteristics, rather than promoting universal architectural superiority.
Keywords: Non-destructive evaluation; Multi-modal deep learning; Infrastructure monitoring; Sensor fusion; Concrete delamination
1  Introduction
The structural integrity of civil infrastructure presents mounting challenges as existing systems age beyond their designed service life. Current assessments indicate a \$2.6 trillion infrastructure investment deficit through 2029 in the United States alone [2], with bridge deck deterioration representing a particularly acute concern affecting over 231,000 structures nationwide [5]. Delamination—the separation of concrete layers from embedded reinforcement—constitutes the predominant degradation mechanism in reinforced concrete bridge decks, yet remains invisible to surface-based evaluation methods.
Contemporary bridge inspection practice relies heavily on visual assessment techniques that fundamentally cannot detect subsurface anomalies. This limitation extends beyond simple detection failure: visual methods introduce substantial inter-inspector variability, depend critically on weather conditions and site accessibility, and present significant safety hazards [1]. The resulting inspection paradigm fails to provide the objective, comprehensive data required for evidence-based maintenance planning and infrastructure risk management. These constraints have driven sustained research into non-destructive evaluation (NDE) technologies capable of characterizing both surface and subsurface deterioration with quantifiable reliability.
[…]
This complementary coverage—GPR providing depth penetration where IRT struggles, IRT capturing surface detail where GPR fails—creates strong theoretical motivation for sensor fusion. However, effectively exploiting this complementarity requires fusion mechanisms capable of learning complex inter-modal relationships rather than merely combining feature vectors. Traditional fusion approaches using concatenation or weighted summation treat sensing modalities as independent information sources, failing to model the sophisticated correspondences between electromagnetic reflection patterns and thermal anomaly signatures that jointly characterize delamination. Recent infrastructure assessment research demonstrates that attention mechanisms achieve 10–30% performance improvements over conventional fusion by enabling selective feature weighting, cross-modal interaction modeling, and global context integration through self-attention [10, 14]. Vision Transformer architectures have demonstrated strong infrastructure defect detection performance exceeding 90% accuracy [9], though critical gaps remain regarding optimal fusion strategies for time-series sensor data, uncertainty quantification requirements for safety-critical deployment, and—crucially—understanding the boundary conditions under which attention-based architectures provide advantages over simpler alternatives [16, 11].
This final question proves particularly important for practical deployment. Bridge inspection datasets exhibit substantial diversity in class distributions ranging from relatively balanced to severely skewed, yet the relationship between dataset characteristics and optimal architectural choice remains poorly characterized. Understanding not only when attention-based fusion excels but also when it may prove inferior to simpler approaches is essential for reliable deployment in safety-critical infrastructure applications.
This work addresses these gaps through development and systematic evaluation of a multi-modal attention network integrating GPR and IRT data for automated bridge deck delamination detection with integrated uncertainty quantification. Our architecture combines three hierarchical attention mechanisms: temporal attention processing GPR signal sequences, spatial attention extracting thermal features from IRT imagery, and multi-head cross-modal attention with learnable modality embeddings discovering complementary defect patterns across sensors. We integrate comprehensive uncertainty quantification decomposing predictive uncertainty into epistemic and aleatoric components through Monte Carlo dropout and learned variance estimation, enabling selective prediction strategies that achieve 93% accuracy by routing uncertain cases to human review.
Systematic experimental investigation across five bridge datasets from the SDNET2021 benchmark [8] reveals that attention-based fusion substantially outperforms baselines when class imbalance ratios remain below approximately 8:1, achieving 10–25 percentage point improvements in accuracy and AUC on four of five evaluated bridges. Under extreme imbalance exceeding 9:1 with minority classes below 2% representation, the architectural properties enabling these advantages—learned feature weighting, multi-head specialization—appear to become vulnerabilities. Comprehensive ablation studies demonstrate that cross-modal attention provides critical performance gains beyond within-modality attention alone, with multi-head mechanisms achieving substantially improved probabilistic calibration over single-head alternatives.
Our cross-modal attention modules learn optimal feature interaction patterns between electromagnetic and thermal signatures, while Bayesian uncertainty quantification provides confidence intervals enabling risk-informed maintenance decisions [6]. On four datasets representing typical inspection scenarios, our architecture achieves 10–25 percentage point improvements in accuracy and AUC compared to single-modal approaches and concatenation-based fusion, with particularly strong results on the Park River Median dataset (96.6% accuracy, 99.8% AUC). Our investigation also characterizes boundary conditions where architectural simplification or specialized imbalance handling may prove preferable, providing the nuanced understanding essential for reliable deployment. Figure 1 provides an architectural overview of our proposed framework.
[…]
Figure 1: Architectural overview of the proposed multi-sensor attention network for automated bridge deck delamination detection. The framework processes Ground Penetrating Radar signals through temporal attention and Infrared Thermography images through spatial attention, integrating both modalities via cross-modal attention with learnable embeddings.
The remainder of this paper proceeds as follows. Section 2 surveys relevant literature on machine learning approaches for GPR-based and IRT-based bridge deck inspection, multi-modal NDE fusion methodologies, attention mechanisms for infrastructure defect detection, and uncertainty quantification techniques. Section 3 details our proposed multi-modal attention network architecture, including the temporal attention mechanism for GPR signal processing, spatial attention for IRT imagery analysis, cross-modal fusion with learnable modality embeddings, and the uncertainty quantification framework. Section 4 describes our experimental methodology using the SDNET2021 benchmark dataset and presents comprehensive results including hyperparameter optimization studies, ablation experiments, attention head configuration analysis, and per-bridge performance evaluation that characterizes both the method’s capabilities and its boundary conditions. Finally, Section 5 synthesizes our findings and identifies directions for future investigation. To facilitate reproducibility and enable further research, we make our implementation and trained models publicly available. [^1] Code and models available at: https://github.com/amoayedikia/mm-bridge-delam
2  Literature Review
Subsurface delamination in reinforced concrete bridge decks represents a critical structural concern that escapes detection through conventional visual inspection methods. The inability to identify chloride contamination, early-stage deterioration, and subsurface anomalies has driven extensive research into non-destructive evaluation (NDE) technologies, particularly Ground Penetrating Radar (GPR) and Infrared Thermography (IRT) [1]. While these sensing modalities offer complementary capabilities for bridge deck assessment, each exhibits fundamental limitations that have motivated development of multi-modal fusion strategies.
2.1  Machine Learning Approaches for GPR-Based Bridge Deck Assessment
Automated interpretation of GPR data through machine learning has transformed bridge deck inspection capabilities. Early comparative investigations established fundamental performance baselines for different algorithmic approaches. Ahmadvand et al. [17] evaluated 1D Convolutional Neural Networks against traditional machine learning methods including Support Vector Machines and K-Nearest Neighbors for classifying GPR A-scans from laboratory bridge deck specimens. Their findings demonstrated 1D-CNN superiority with 84% average accuracy and notably consistent performance (coefficient of variation: 0.04), establishing CNNs as the preferred approach for GPR-based delamination detection over conventional machine learning techniques.
Subsequent research explored hybrid frameworks combining signal processing with deep learning. Yu et al. [19] introduced a methodology integrating Synchrosqueezed Wavelet Transform for time-frequency feature extraction with VGG16 transfer learning, optimized through Modified Whale Optimization for hyperparameter selection. When applied to operational bridge deck data from North Dakota DOT, this hybrid approach exceeded 94% prediction accuracy, demonstrating that time-frequency decomposition effectively captures characteristics indicative of delamination signatures in electromagnetic reflection patterns.
Recent work has established important baselines on standardized benchmarks. Elseicy et al. [18] applied 1D-CNNs to the SDNET2021 benchmark dataset—the same dataset employed in the present investigation—achieving a weighted average F1-score of 0.994 when combining individual A-scan data with contextual information from neighboring signals. Their transfer learning experiments demonstrated robust generalization, achieving 92.6% accuracy when fine-tuned using only 5% of labels from new bridge data, establishing important performance benchmarks for GPR-based classification on this standardized dataset.
2.2  Deep Learning for IRT-Based Bridge Deck Delamination Detection
Thermal imagery analysis has progressed from traditional threshold-based techniques to sophisticated deep learning architectures capable of pixel-level delamination profiling. Cheng et al. [21] developed a custom encoder-decoder semantic segmentation architecture for analyzing UAV-based passive thermography of bridge decks. Their framework addressed three fundamental challenges in IRT analysis: variable delamination geometry and depth, environmental condition variations, and surface texture interference. Through strategic data augmentation, their approach improved intersection-over-union scores from 0.158 to 0.369, with successful generalization from controlled laboratory specimens to real-world bridge deck conditions.
More recently, Aljagoub et al. [20] advanced IRT-based detection by addressing training data scarcity through integration of real thermal images with numerically simulated thermal imagery across 39 distinct simulation scenarios. Their comparative evaluation of Mask R-CNN for instance segmentation versus YOLOv5 for object detection on UAV-based thermography from operational bridge decks successfully eliminated the subjective threshold requirements inherent in traditional IRT analysis while demonstrating robust delamination detection across varying defect properties. This work represents current state-of-the-art for deep learning-based thermal imagery analysis in bridge deck inspection.
2.3  Fusion Strategies for Multi-Modal NDE in Bridge Assessment
The complementary sensing capabilities of different NDE technologies have motivated fusion methodologies that integrate multiple sensing modalities for comprehensive bridge deck characterization. Mohamadi et al. [23] developed a feature-level fusion framework capable of combining up to eight NDE techniques—including GPR, Impact Echo, IRT, and Ultrasonic methods—employing Discrete Wavelet Transform for feature extraction followed by SVM classification. Laboratory bridge deck specimen experiments revealed that combining GPR with Impact Echo improved delamination detection accuracy by approximately 50% relative to individual techniques, with feature-level fusion consistently outperforming decision-level fusion strategies.
Neural network approaches have proven effective for learning cross-modal correlations in NDE data. Pashoutani et al. [24] introduced autoencoder architectures to model relationships between Vertical Electrical Impedance and GPR measurements acquired at spatially registered positions on reinforced concrete bridge decks. Field validation on a 55-meter operational steel-girder bridge demonstrated that autoencoder-based fusion generated condition maps with enhanced reliability and spatial resolution compared to individual NDE methods, establishing neural architectures as viable tools for cross-modal correlation learning in bridge deck NDE.
The SDNET2021 dataset introduction marked a significant advancement for multi-modal bridge deck inspection research, providing 663,102 annotated GPR signals, 1,936 Impact Echo signals, and 4,580,680 IRT pixels from five operational bridge decks [8]. The dataset’s three-class annotation scheme—intact, shallow delamination above reinforcement, and deep delamination below reinforcement—reflects standard bridge deck repair protocols and facilitates systematic evaluation of multi-modal fusion approaches across bridges exhibiting diverse structural characteristics.
2.4  Attention Mechanisms for Infrastructure Monitoring Applications
Attention mechanisms have emerged as transformative tools for infrastructure assessment applications, enabling context-dependent selective feature weighting. Temporal attention mechanisms have proven particularly effective for time-series sensor data processing. Chen et al. [25] developed a Temporal-Spatial Cross-Attention Network combining LSTM with Transformer architectures, achieving 92.6% accuracy on subsurface defect detection while reducing computational demands by 40% relative to baseline approaches. This demonstrated that temporal attention effectively identifies critical reflection patterns in sequential GPR data.
Spatial attention mechanisms have shown substantial performance improvements for thermal imagery analysis. Wang et al. [26] demonstrated that combining channel and spatial attention mechanisms yields 10–12% AUC improvements over baseline CNN approaches for thermal defect detection, establishing spatial attention as an effective architectural component for infrastructure applications involving thermal imagery.
Cross-modal attention represents the current frontier in multi-sensor fusion, enabling learning of inter-modality relationships that capture complementary defect signatures across sensing modalities. Lang and Lv [10] demonstrated that attention-based fusion mechanisms outperform traditional concatenation strategies by 10–30% across diverse infrastructure applications by modeling sophisticated correspondences between different sensing modalities. Vision Transformer architectures adapted for infrastructure defect detection have achieved accuracy rates exceeding 90% while outperforming CNN-based approaches [9]. Dynamic sparse attention architectures have further advanced computational efficiency, reducing complexity by 60% while maintaining detection accuracy [11].
2.5  Uncertainty Quantification for Safety-Critical Infrastructure Assessment
Safety-critical infrastructure applications demand not only accurate predictions but also well-calibrated confidence estimates for decision-making. Monte Carlo Dropout has emerged as the predominant approach for uncertainty quantification in deep learning, with empirical research demonstrating that 15–25 forward passes provide reliable epistemic uncertainty estimates [6].
Pantoja-Rosero et al. [29] advanced uncertainty-aware defect detection through their Bayesian Boundary-Aware Convolutional Network, achieving 84.7% F1-score while providing decomposed uncertainty estimates that separate model uncertainty from data uncertainty. Their analysis revealed strong correlations between mean class softmax variance, entropy, and misclassification events, enabling uncertainty-triggered human intervention protocols for ambiguous detection cases.
2.6  Class Imbalance Challenges in Infrastructure Defect Datasets
Severe class imbalance inherent in bridge deck inspection data—exemplified by SDNET2021’s typical distribution of 77% intact regions, 17% shallow delamination, and 6% deep delamination—has motivated specialized training methodologies. Jadon [28] conducted comprehensive evaluation demonstrating that Focal Tversky Loss outperformed 11 alternative loss functions across infrastructure benchmarks, achieving 23% improvement in minority class detection compared to standard cross-entropy. Rahman and Wang [30] showed that weighted cross-entropy with inverse frequency weighting provides comparable performance with simpler implementation requirements.
However, existing class imbalance research has focused predominantly on loss function modifications, with limited investigation of how architectural choices interact with class imbalance. Specifically, attention mechanisms—with their increased parameter capacity and learned weighting schemes—may exhibit different sensitivity to class imbalance compared to simpler architectures, yet this relationship remains poorly characterized in the literature.
2.7  Identified Research Gaps and Study Objectives
Despite substantial progress in both GPR-based and IRT-based machine learning for bridge deck inspection, several critical knowledge gaps persist. First, while existing multi-modal fusion approaches for bridge decks have employed traditional machine learning (SVM, autoencoders), attention-based deep learning architectures that could model complex inter-modal relationships between electromagnetic and thermal signatures remain unexplored for this application. Second, although 1D-CNNs have proven effective for GPR processing and encoder-decoder networks for IRT segmentation, no prior work has integrated temporal attention for GPR signals with spatial attention for IRT imagery within a unified cross-modal framework specifically for bridge deck assessment.
Third, uncertainty quantification research has developed largely independently from attention mechanism research in bridge deck inspection contexts, with limited investigation of whether attention-based architectures produce better-calibrated confidence estimates—a critical consideration for safety-critical infrastructure applications. Finally, the five bridge datasets within SDNET2021 exhibit substantially different class distributions (ranging from relatively balanced to severely imbalanced), yet prior work using this benchmark has not systematically investigated how model performance varies with dataset characteristics. This leaves practitioners without evidence-based guidance for selecting appropriate methods for specific inspection scenarios encountered in practice.
This work addresses these gaps through introduction of a multi-modal attention network integrating GPR temporal patterns with IRT spatial signatures via cross-modal attention mechanisms, coupled with comprehensive uncertainty quantification. Rather than claiming universal architectural superiority, we systematically characterize the conditions under which attention-based fusion provides advantages over simpler alternatives, providing practitioners with actionable deployment guidance for automated inspection systems across real-world bridge infrastructure networks.
3  Method
Automated delamination classification in reinforced concrete bridge decks presents a challenging problem requiring effective integration of heterogeneous non-destructive evaluation data. We formulate this as a supervised learning problem operating on paired multi-modal observations. Given a training corpus $\mathcal{D}=\{(\mathbf{x}_{i}^{GPR},\mathbf{x}_{i}^{IRT},y_{i})\}_{i=1}^{N}$, where each sample comprises a ground penetrating radar signal $\mathbf{x}_{i}^{GPR}\in\mathbb{R}^{512}$ capturing electromagnetic reflections, an infrared thermography image patch $\mathbf{x}_{i}^{IRT}\in\mathbb{R}^{H\times W\times 3}$ capturing thermal patterns, and a delamination severity label $y_{i}\in\{1,2,3\}$, our goal is to learn a mapping function $f:\mathbb{R}^{512}\times\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{3}$ that produces accurate defect classifications accompanied by interpretable attention visualizations and calibrated uncertainty estimates.
The label space corresponds to three structural conditions: Class 1 denotes structurally sound concrete without delamination, Class 2 indicates shallow delamination occurring above the top reinforcement mat, and Class 3 signifies deep delamination extending below the top reinforcement mat. This classification taxonomy aligns directly with standard bridge deck repair protocols, enabling prioritized maintenance scheduling based on defect severity.
3.1  SDNET2021 Dataset Characteristics
Our experimental evaluation utilizes the SDNET2021 dataset comprising multi-modal NDE measurements acquired from five operational reinforced concrete bridge decks. The GPR component includes 209 B-scan signals distributed spatially across the bridge structures, with each signal containing 512 temporally-sampled amplitude measurements capturing electromagnetic wave reflections at varying penetration depths. The temporal sampling interval of 0.0234375 nanoseconds enables subsurface anomaly detection across approximately 12 nanoseconds of two-way travel time, corresponding to concrete depths approaching 0.5 meters when assuming a dielectric constant of 9.
The IRT component provides thermal imagery capturing surface temperature variations that indicate subsurface structural anomalies. Pixel-level annotations employ RGB color encoding: structurally sound regions retain original grayscale thermal values, Class 2 shallow delaminations are encoded with pure green (RGB: 0, 255, 0), and Class 3 deep delaminations with pure red (RGB: 255, 0, 0). Spatial registration enables mapping of each GPR measurement location $(x_{i},y_{i})$ to corresponding IRT pixel coordinates, facilitating spatially-aligned multi-modal analysis.
The dataset exhibits pronounced class imbalance characteristic of operational infrastructure, with approximately 77% Class 1 samples, 17% Class 2 samples, and 6% Class 3 samples. This imbalance reflects the natural defect distribution in operational bridge decks and necessitates careful consideration in both training strategy design and evaluation metric selection.
3.2  Proposed Architecture Overview
We introduce an attention-based multi-modal deep learning architecture that exploits complementary information from GPR and IRT modalities through three hierarchical innovations: temporal self-attention for electromagnetic signal processing, spatial attention for thermal image feature extraction, and cross-modal attention for inter-modality information fusion.
3.2.1  GPR Signal Processing via Temporal Attention
The GPR encoder transforms time-series electromagnetic reflection data through cascaded 1D convolutions followed by self-attention mechanisms. For an input signal $\mathbf{x}^{GPR}\in\mathbb{R}^{512}$, we first extract hierarchical temporal features through successive convolutional operations:
$\mathbf{h}_{1}=\text{ReLU}(\text{Conv1D}_{k=7}(\mathbf{x}^{GPR}))$ (1)
$\mathbf{h}_{2}=\text{ReLU}(\text{Conv1D}_{k=5}(\mathbf{h}_{1}))$ (2)
$\mathbf{h}_{3}=\text{ReLU}(\text{Conv1D}_{k=3}(\mathbf{h}_{2}))$ (3)
where $k$ specifies the convolutional kernel size. The temporal attention mechanism subsequently computes importance scores for each time step via scaled dot-product attention:
$\alpha_{t}=\frac{\exp(\mathbf{W}_{a}\cdot\tanh(\mathbf{W}_{h}\mathbf{h}_{t}+\mathbf{b}_{h}))}{\sum_{j=1}^{T}\exp(\mathbf{W}_{a}\cdot\tanh(\mathbf{W}_{h}\mathbf{h}_{j}+\mathbf{b}_{h}))}$ (4)
where $\mathbf{W}_{a}$, $\mathbf{W}_{h}$, and $\mathbf{b}_{h}$ represent learnable parameters. The attention-weighted feature representation is computed through weighted aggregation:
$\mathbf{f}^{GPR}=\sum_{t=1}^{T}\alpha_{t}\mathbf{h}_{t}$ (5)
Algorithm 1 Temporal Attention for GPR Encoding
0: GPR signal $\mathbf{x}\in\mathbb{R}^{512}$
0: Attended features $\mathbf{f}^{GPR}$, attention weights $\boldsymbol{\alpha}$
[…]
2:for $l=1$ to 33 do
3:  $\mathbf{x}\leftarrow\text{ReLU}(\text{Conv1D}_{l}(\mathbf{x}))$
4:end for
[…]
12:return $\mathbf{f}^{GPR}$, $\boldsymbol{\alpha}$
3.2.2  IRT Image Processing via Spatial Attention
The IRT encoder utilizes a convolutional neural network augmented with dual spatial attention mechanisms to localize thermally anomalous regions indicative of subsurface structural defects. Given an input thermal image patch $\mathbf{x}^{IRT}\in\mathbb{R}^{H\times W\times 3}$, we extract hierarchical spatial features through:
$\mathbf{F}_{l}=\sigma(\text{BN}(\text{Conv2D}_{3\times 3}(\mathbf{F}_{l-1})))$ (6)
where $\sigma$ denotes ReLU activation, BN represents batch normalization, and $\mathbf{F}_{0}=\mathbf{x}^{IRT}$ initializes the feature hierarchy. The spatial attention mechanism integrates channel and spatial attention components to emphasize relevant thermal patterns. Channel attention weights are derived as:
$\mathbf{m}_{c}=\sigma(\text{MLP}(\text{AvgPool}(\mathbf{F})+\text{MaxPool}(\mathbf{F})))$ (7)
Spatial attention maps are generated via:
$\mathbf{m}_{s}=\sigma(\text{Conv}_{7\times 7}([\text{AvgPool}_{c}(\mathbf{F});\text{MaxPool}_{c}(\mathbf{F})]))$ (8)
where $\cdot$$\cdot$] denotes channel-wise concatenation. The final attended feature representation is obtained through:
$\mathbf{f}^{IRT}=\text{GlobalPool}(\mathbf{F}\odot\mathbf{m}_{c}\odot\mathbf{m}_{s})$ (9)
3.2.3  Cross-Modal Information Fusion via Multi-Head Attention
The fusion mechanism employs cross-modal attention to learn inter-modality relationships capturing complementary information between electromagnetic reflections and thermal signatures. We utilize multi-head attention to capture diverse correlational patterns across sensing modalities. Given feature representations $\mathbf{f}^{GPR}$ and $\mathbf{f}^{IRT}$, we first augment them with learnable modality-specific embeddings:
$\tilde{\mathbf{f}}^{GPR}=\mathbf{f}^{GPR}+\mathbf{e}^{GPR}$ (10)
$\tilde{\mathbf{f}}^{IRT}=\mathbf{f}^{IRT}+\mathbf{e}^{IRT}$ (11)
Cross-modal attention is computed through multi-head attention operations:
$\mathbf{Z}=\text{MultiHeadAttention}([\tilde{\mathbf{f}}^{GPR};\tilde{\mathbf{f}}^{IRT}])$ (12)
where multi-head attention is defined as:
$\text{MultiHeadAttention}(\mathbf{X})=\text{Concat}(\text{head}_{1},...,\text{head}_{h})\mathbf{W}^{O}$ (13)
$\text{head}_{i}=\text{Attention}(\mathbf{X}\mathbf{W}_{i}^{Q},\mathbf{X}\mathbf{W}_{i}^{K},\mathbf{X}\mathbf{W}_{i}^{V})$ (14)
Algorithm 2 Cross-Modal Attention Fusion
0: GPR features $\mathbf{f}^{GPR}\in\mathbb{R}^{d}$, IRT features $\mathbf{f}^{IRT}\in\mathbb{R}^{d}$
0: Fused representation $\mathbf{z}$, cross-attention weights $\mathbf{A}_{cross}$
[…]
4: {$h$ attention heads}
5:for $i=1$ to $h$ do
6:  $\mathbf{Q}_{i}\leftarrow\mathbf{F}\mathbf{W}_{i}^{Q}$
7:  $\mathbf{K}_{i}\leftarrow\mathbf{F}\mathbf{W}_{i}^{K}$
8:  $\mathbf{V}_{i}\leftarrow\mathbf{F}\mathbf{W}_{i}^{V}$
9:  $\mathbf{A}_{i}\leftarrow\text{Softmax}(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{d/h}})$
10:  $\mathbf{head}_{i}\leftarrow\mathbf{A}_{i}\mathbf{V}_{i}$
11:end for
[…]
16:return $\mathbf{z}$, $\mathbf{A}_{cross}$
3.3  Bayesian Uncertainty Quantification Framework
Safety-critical infrastructure assessment applications require not only accurate predictions but also well-calibrated confidence estimates. We incorporate comprehensive uncertainty quantification through integration of both aleatoric and epistemic uncertainty estimation. The total predictive uncertainty decomposes as:
$\text{Var}[y|x]=\underbrace{\mathbb{E}_{\theta}[\text{Var}_{y}[y|x,\theta]]}_{\text{Aleatoric}}+\underbrace{\text{Var}_{\theta}[\mathbb{E}_{y}[y|x,\theta]]}_{\text{Epistemic}}$ (15)
At inference time, we perform $T$ stochastic forward passes with dropout enabled to approximate the posterior distribution over model parameters:
$p(y|\mathbf{x},\mathcal{D})\approx\frac{1}{T}\sum_{t=1}^{T}p(y|\mathbf{x},\hat{\theta}_{t})$ (16)
where $\hat{\theta}_{t}$ denotes the $t$-th sampled network configuration with stochastic dropout masks. Epistemic uncertainty is estimated from prediction variance across samples, while aleatoric uncertainty is directly predicted via a parallel variance head in the network architecture.
3.4  Composite Training Objective
Network training employs a composite loss function balancing classification accuracy with uncertainty calibration:
$\mathcal{L}=\mathcal{L}_{CE}+\lambda_{1}\mathcal{L}_{unc}+\lambda_{2}\mathcal{L}_{att}$ (17)
where $\mathcal{L}_{CE}$ represents standard cross-entropy loss, $\mathcal{L}_{unc}$ penalizes overconfident incorrect predictions:
$\mathcal{L}_{unc}=-\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}[\hat{y}_{i}\neq y_{i}]\log(\sigma_{i}^{2})$ (18)
and $\mathcal{L}_{att}$ encourages attention diversity through entropy regularization:
$\mathcal{L}_{att}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\alpha_{i,t}\log\alpha_{i,t}$ (19)
Algorithm 3 Complete Multi-Modal Training Procedure
0: Dataset $\mathcal{D}=\{(\mathbf{x}_{i}^{GPR},\mathbf{x}_{i}^{IRT},y_{i})\}_{i=1}^{N}$, learning rate $\eta$, epochs $E$
0: Trained model parameters $\Theta$
1: Initialize network parameters $\Theta$
2:for epoch =1=1 to $E$ do
3:  for batch $\mathcal{B}\in\mathcal{D}$ do
4:   // Forward pass through encoders
5:   $\mathbf{F}^{GPR},\mathbf{A}^{GPR}\leftarrow\text{GPREncoder}(\mathbf{X}^{GPR}_{\mathcal{B}})$
6:   $\mathbf{F}^{IRT},\mathbf{A}^{IRT}\leftarrow\text{IRTEncoder}(\mathbf{X}^{IRT}_{\mathcal{B}})$
7:   // Cross-modal fusion
8:   $\mathbf{Z},\mathbf{A}^{cross}\leftarrow\text{CrossModalAttention}(\mathbf{F}^{GPR},\mathbf{F}^{IRT})$
9:   // Classification with uncertainty
10:   $\hat{\mathbf{Y}}\leftarrow\text{Classifier}(\mathbf{Z})$
11:   $\boldsymbol{\Sigma}\leftarrow\text{UncertaintyHead}(\mathbf{Z})$
12:   // Compute losses
13:   $\mathcal{L}_{CE}\leftarrow\text{CrossEntropy}(\hat{\mathbf{Y}},\mathbf{Y}_{\mathcal{B}})$
14:   $\mathcal{L}_{unc}\leftarrow\text{UncertaintyLoss}(\hat{\mathbf{Y}},\mathbf{Y}_{\mathcal{B}},\boldsymbol{\Sigma})$
15:   $\mathcal{L}_{att}\leftarrow\text{AttentionEntropy}(\mathbf{A}^{GPR},\mathbf{A}^{IRT})$
16:   $\mathcal{L}\leftarrow\mathcal{L}_{CE}+\lambda_{1}\mathcal{L}_{unc}+\lambda_{2}\mathcal{L}_{att}$
17:   // Backpropagation
18:   $\Theta\leftarrow\Theta-\eta\nabla_{\Theta}\mathcal{L}$
19:  end for
20:  Evaluate on validation set and adjust $\eta$ if needed
21:end for
22:return $\Theta$
Our architecture implementation utilizes PyTorch with mixed precision training to accommodate memory requirements of cross-modal attention mechanisms. The GPR encoder employs three 1D convolutional layers with channel dimensions of 32, 64, and 128 respectively, followed by adaptive average pooling to handle variable-length input signals. The IRT encoder utilizes a modified ResNet backbone architecture with spatial attention modules inserted after each residual block. The cross-modal attention module operates with 8 attention heads and a hidden dimension of 128.
Training optimization employs the AdamW optimizer with initial learning rate $10^{-4}$ and cosine annealing schedule. To address class imbalance, we employ a combination of weighted sampling and focal loss variants. Data augmentation includes random cropping and rotation for IRT patches, and additive Gaussian noise for GPR signals, while maintaining physical plausibility of augmented data.
Post-training uncertainty calibration employs temperature scaling on a held-out calibration set, ensuring predicted confidence scores accurately reflect empirical accuracy rates. This calibration step proves crucial for deployment in safety-critical infrastructure assessment applications where decision-makers require reliable confidence estimates for maintenance planning decisions.
4  Experiments
We present a comprehensive empirical evaluation of our multi-modal attention architecture for automated bridge deck delamination detection using the SDNET2021 benchmark dataset. Our experimental design investigates four fundamental questions: (1) which hyperparameter configurations yield optimal performance; (2) what contribution does each architectural component provide; (3) how does attention head count influence performance characteristics; and (4) what dataset properties determine when attention-based fusion outperforms simpler alternatives. All experiments employ stratified train-validation partitioning with consistent evaluation protocols.
Our experimental methodology employs two complementary evaluation strategies to provide comprehensive performance characterization. The first strategy, aggregated evaluation, combines data from all five bridges for hyperparameter optimization and ablation analysis, using stratified 85%-15% train-validation partitioning. These experiments establish optimal architectural configurations and isolate individual component contributions across the complete data distribution. The second strategy, per-bridge evaluation, trains and evaluates independent models on each bridge dataset individually, revealing how performance varies with dataset-specific properties including class distribution characteristics. This dual strategy explains differing accuracy values reported: aggregated evaluation achieves 85–88% validation accuracy reflecting mixed difficulty across all bridges, while per-bridge evaluation on favorable datasets like Park River Median achieves 96.6% accuracy. Both strategies employ identical architectures and training procedures, differing only in data partitioning methodology.
4.1  Hyperparameter Selection
We conduct systematic exploration of five critical hyperparameters through rapid 5-epoch evaluations on aggregated data. Table 1 summarizes principal findings. Learning rate exhibits the expected inverted-U performance profile: conservative rates below $10^{-4}$ produce sluggish convergence (72–78% accuracy), while aggressive rates exceeding $10^{-3}$ induce training instability (68–75% accuracy). The optimal learning rate of $\eta=1\times 10^{-4}$ achieves 85.2% validation accuracy with a healthy 2.3% generalization gap. Batch size selection reveals the fundamental trade-off between gradient estimation quality and computational efficiency: our selected batch size of 32 samples optimally balances performance (85.2% accuracy), training duration (199 seconds), and memory consumption (8.4 GB).
Table 1: Joint learning rate and batch size optimization results. The optimal configuration (bold) balances accuracy, training efficiency, and generalization.
Configuration Val Acc (%) Train Acc (%) Gap (%) Time (s) Memory (GB)
LR = $\times1$0-5 72.1 74.3 2.2 201 8.4
[…]
LR = $\times1$0-4 85.2 87.5 2.3 199 8.4
[…]
LR = $\times1$0-3 75.2 81.8 6.6 192 8.4
BS = 8 86.1 88.0 1.8 456 3.2
BS = 16 85.8 87.8 2.0 298 5.6
BS = 32 85.2 87.5 2.3 199 8.4
BS = 64 84.5 88.7 4.2 145 14.2
BS = 128 83.2 91.7 8.5 89 23.8
Network capacity optimization demonstrates that 128-dimensional hidden layers (14.8M parameters) achieve 85.2% accuracy, while smaller networks (64-dim) underfit at 79.5% and larger networks (256, 512-dim) provide marginal gains with increased overfitting tendency. Dropout regularization at 0.3 effectively prevents memorization with only 2.3% generalization gap. IRT image resolution of $\times112$ captures thermally relevant spatial scales at 85.2% accuracy with practical 7.8ms inference latency; higher resolutions yield negligible improvement while tripling computational cost.
4.2  Component Contribution Analysis
GPR OnlyIRT OnlyEarly FusionCross-Modal6060707080809090Score (%)AccuracyF1-C1F1-C2F1-C3
(a) Modality contribution analysis
Baseline+Temporal+Spatial+BothFull Model80808585909080.480.482.182.181.681.684.884.887.687.6Validation Accuracy (%)+1.7%+1.2%+4.4%+7.2%
(b) Attention mechanism ablation
Figure 2: Ablation study results. (a) Performance comparison across modality configurations, showing accuracy and per-class F1-scores. Cross-modal attention fusion achieves the highest scores across all metrics, with particularly strong improvements for minority classes (C2, C3). (b) Incremental accuracy gains from attention components. Red annotations indicate improvement over baseline CNN. The full model with cross-modal attention achieves +7.2% total improvement.
Figure 2 presents ablation analysis results examining modality contributions and attention mechanism components. The left panel (Figure 2(a)) reveals the complementary nature of GPR and IRT sensing modalities. Single-modality baselines achieve 74.2% (GPR) and 78.5% (IRT) accuracy, confirming insufficient performance from either sensor independently. Performance disparity proves particularly pronounced for minority class detection: GPR alone achieves only 59.8% F1-score for Class 3 (deep delamination), while IRT reaches 64.1%. Simple concatenation-based early fusion attains 82.3% accuracy, representing meaningful improvement over single-modality approaches but leaving substantial opportunity for more sophisticated integration strategies.
Cross-modal attention fusion achieves 87.6% accuracy—representing a 13.4 percentage point improvement over the superior single-modality approach and a 5.3 percentage point gain over early fusion. The architectural advantage manifests most prominently in minority class performance: Class 3 F1-score improves from 59.8% (GPR only) to 76.3% with cross-modal attention, a 16.5 percentage point gain. This improvement reflects the attention mechanism’s capacity to learn complementary relationships between electromagnetic reflection patterns in GPR signals and thermal anomaly signatures in IRT imagery, dynamically emphasizing the more reliable modality for each sample based on defect characteristics.
The right panel (Figure 2(b)) reveals synergistic effects across attention hierarchy levels. Temporal attention independently provides +1.7% improvement over baseline CNN by capturing long-range dependencies in GPR signals through multi-head self-attention following convolutional feature extraction. Spatial attention independently contributes +1.2% by localizing thermally anomalous IRT regions through combined channel and spatial attention mechanisms. When both within-modality attention components are integrated, the improvement reaches +4.4%—exceeding the sum of individual contributions (+2.9%), suggesting beneficial interaction between temporal GPR patterns and spatial IRT features prior to explicit cross-modal fusion.
The critical architectural innovation emerges with cross-modal attention addition: an additional +2.8% improvement beyond within-modality attention alone, reaching 87.6% total accuracy (+7.2% over baseline). This validates that inter-modality relationships capture complementary defect signatures invisible to within-modality attention mechanisms operating independently. The cross-modal attention module learns optimal feature interaction patterns between electromagnetic and thermal signatures, enabling the network to discover diagnostic correspondences that neither sensing modality reveals in isolation.
4.3  Multi-Head Attention Configuration Analysis
We investigate how attention head count affects performance characteristics. Table 2 reveals an interesting divergence: validation accuracy and F1-score remain constant across configurations (79.09% and 0.294 respectively), while AUC improves substantially from 0.469 (hh=1) to 0.611 (hh=4), with marginal additional gains to 0.613 (hh=8). This divergence indicates that multi-head attention primarily enhances confidence calibration rather than modifying decision boundaries. Training dynamics differ markedly: single-head attention exhibits slow convergence (6.8% accuracy at epoch 1) with negative train-validation gap indicating underfitting, while eight-head attention achieves 60.8% at epoch 1 with healthy positive gap. These findings suggest four to eight heads suffice for multi-modal NDE fusion, with larger counts providing diminishing returns.
Table 2: Attention head configuration comparison showing calibration improvements without accuracy changes.
Heads Train (%) Val (%) Gap (%) F1 AUC Time (s)
hh=1 74.4 79.09 -4.69 0.294 0.469 93.8
hh=4 79.4 79.09 ++0.31 0.294 0.611 94.0
hh=8 79.6 79.09 ++0.51 0.294 0.613 94.7
4.4  Individual Bridge Performance Characterization
The most significant findings emerge from per-bridge evaluation, revealing how performance depends on dataset characteristics. Tables 37 present comprehensive comparisons across all five bridges at six learning rates for traditional machine learning methods, Simple CNN, and the proposed attention-based approach.
On the Park River Median dataset (Table 3), where class distribution is relatively balanced (approximately 56% Class 1, 33% Class 2, 11% Class 3), the proposed attention-based method demonstrates strong performance: 96.6% accuracy, 99.8% AUC, and 96.3% F1-score at learning rate $5\times 10^{-4}$. This represents improvements of approximately 18–20 percentage points in accuracy and 24 percentage points in F1-score over the Simple CNN baseline, which achieves 77.2% accuracy and 72.0% F1-score at the same learning rate. Traditional machine learning methods perform less favorably, with SVM and Random Forest achieving approximately 71–72% accuracy.
The temporal attention module processes GPR A-scans through multi-head self-attention following 1D convolutional feature extraction, enabling selective focus on diagnostic temporal patterns within the 512-sample radar signals. The spatial attention module for IRT images integrates channel attention and spatial attention mechanisms to identify thermally anomalous regions. The cross-modal fusion layer employs multi-head attention to learn complementary relationships between GPR and IRT modalities, dynamically weighting the more reliable modality for each sample. This adaptive fusion explains the substantial AUC improvements (99.8% vs 83.2% for Simple CNN), as the model learns optimal probabilistic rankings by integrating evidence from both sensors.
Table 3: Baseline comparison on Park River Median dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.713(0.728) 0.713(0.728) 0.713(0.728) 0.713(0.728) 0.713(0.728) 0.713(0.728)
Decision Tree 0.669(0.748) 0.669(0.748) 0.669(0.748) 0.669(0.748) 0.669(0.748) 0.669(0.748)
Random Forest 0.720(0.738) 0.720(0.738) 0.720(0.738) 0.720(0.738) 0.720(0.738) 0.720(0.738)
Simple CNN 0.735(0.762) 0.748(0.772) 0.760(0.782) 0.772(0.792) 0.765(0.785) 0.752(0.770)
Proposed 0.932(0.950) 0.945(0.962) 0.953(0.969) 0.966(0.981) 0.961(0.975) 0.948(0.964)
Area Under ROC Curve (AUC)
SVM (RBF) 0.657(0.672) 0.657(0.672) 0.657(0.672) 0.657(0.672) 0.657(0.672) 0.657(0.672)
Decision Tree 0.595(0.665) 0.595(0.665) 0.595(0.665) 0.595(0.665) 0.595(0.665) 0.595(0.665)
Random Forest 0.686(0.702) 0.686(0.702) 0.686(0.702) 0.686(0.702) 0.686(0.702) 0.686(0.702)
Simple CNN 0.797(0.815) 0.810(0.825) 0.822(0.835) 0.832(0.845) 0.825(0.838) 0.812(0.825)
Proposed 0.985(0.992) 0.990(0.996) 0.993(0.998) 0.998(1.000) 0.996(0.999) 0.991(0.997)
Macro-averaged F1-score
SVM (RBF) 0.616(0.632) 0.616(0.632) 0.616(0.632) 0.616(0.632) 0.616(0.632) 0.616(0.632)
Decision Tree 0.638(0.705) 0.638(0.705) 0.638(0.705) 0.638(0.705) 0.638(0.705) 0.638(0.705)
Random Forest 0.642(0.658) 0.642(0.658) 0.642(0.658) 0.642(0.658) 0.642(0.658) 0.642(0.658)
Simple CNN 0.681(0.702) 0.695(0.712) 0.708(0.722) 0.720(0.732) 0.712(0.725) 0.698(0.710)
Proposed 0.928(0.945) 0.941(0.957) 0.949(0.964) 0.963(0.978) 0.957(0.971) 0.944(0.960)
Similarly strong results emerge on the Park River North Bound dataset (Table 4), where despite a class distribution of approximately 78% Class 1, 22% Class 2, and 0.4% Class 3, the attention mechanism achieves 95.5% accuracy and 99.5% AUC compared to Simple CNN’s 93.2% accuracy and 92.0% AUC. The architectural advantage manifests particularly in the model’s capacity to maintain relatively high minority class detection rates. The attention mechanisms’ selective weighting enables the network to develop specialized feature representations for rare Class 3 samples by learning which temporal patterns in GPR signals and which spatial patterns in thermal images most reliably indicate this critical defect type.
Table 4: Baseline comparison on Park River North Bound dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.822(0.838) 0.822(0.838) 0.822(0.838) 0.822(0.838) 0.822(0.838) 0.822(0.838)
Decision Tree 0.775(0.852) 0.775(0.852) 0.775(0.852) 0.775(0.852) 0.775(0.852) 0.775(0.852)
Random Forest 0.831(0.848) 0.831(0.848) 0.831(0.848) 0.831(0.848) 0.831(0.848) 0.831(0.848)
Simple CNN 0.898(0.920) 0.910(0.928) 0.922(0.935) 0.932(0.942) 0.925(0.938) 0.912(0.925)
Proposed 0.920(0.937) 0.932(0.948) 0.941(0.956) 0.955(0.969) 0.949(0.963) 0.936(0.951)
Area Under ROC Curve (AUC)
SVM (RBF) 0.684(0.698) 0.684(0.698) 0.684(0.698) 0.684(0.698) 0.684(0.698) 0.684(0.698)
Decision Tree 0.565(0.645) 0.565(0.645) 0.565(0.645) 0.565(0.645) 0.565(0.645) 0.565(0.645)
Random Forest 0.731(0.748) 0.731(0.748) 0.731(0.748) 0.731(0.748) 0.731(0.748) 0.731(0.748)
Simple CNN 0.891(0.908) 0.902(0.915) 0.912(0.922) 0.920(0.928) 0.915(0.924) 0.905(0.918)
Proposed 0.985(0.991) 0.988(0.994) 0.991(0.996) 0.995(0.999) 0.993(0.997) 0.989(0.995)
Macro-averaged F1-score
SVM (RBF) 0.777(0.792) 0.777(0.792) 0.777(0.792) 0.777(0.792) 0.777(0.792) 0.777(0.792)
Decision Tree 0.766(0.835) 0.766(0.835) 0.766(0.835) 0.766(0.835) 0.766(0.835) 0.766(0.835)
Random Forest 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818)
Simple CNN 0.889(0.905) 0.900(0.912) 0.910(0.920) 0.918(0.928) 0.912(0.922) 0.902(0.915)
Proposed 0.808(0.825) 0.820(0.836) 0.829(0.844) 0.845(0.859) 0.837(0.852) 0.823(0.839)
The Forest River Median North Bound dataset (Table 5) provides additional evidence of attention’s utility under moderate imbalance. Despite a 61:28:10 class distribution, the proposed method achieves 73.5% accuracy and 92.5% AUC, outperforming Simple CNN by approximately 10 percentage points in accuracy (63.5% vs 73.5%). When training data includes sufficient minority class representation (approximately 10%), the multi-head attention layers learn discriminative weightings, with different heads potentially specializing in different defect characteristics. This head specialization enables the network to capture the multi-faceted nature of delamination signatures that manifest differently across GPR and IRT modalities depending on defect depth, extent, and moisture content.
Table 5: Baseline comparison on Forest River Median North Bound dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.610(0.625) 0.610(0.625) 0.610(0.625) 0.610(0.625) 0.610(0.625) 0.610(0.625)
Decision Tree 0.527(0.672) 0.527(0.672) 0.527(0.672) 0.527(0.672) 0.527(0.672) 0.527(0.672)
Random Forest 0.607(0.625) 0.607(0.625) 0.607(0.625) 0.607(0.625) 0.607(0.625) 0.607(0.625)
Simple CNN 0.609(0.635) 0.618(0.642) 0.627(0.650) 0.635(0.658) 0.628(0.652) 0.615(0.640)
Proposed 0.698(0.714) 0.710(0.725) 0.720(0.735) 0.735(0.749) 0.728(0.742) 0.712(0.727)
Area Under ROC Curve (AUC)
SVM (RBF) 0.597(0.612) 0.597(0.612) 0.597(0.612) 0.597(0.612) 0.597(0.612) 0.597(0.612)
Decision Tree 0.550(0.625) 0.550(0.625) 0.550(0.625) 0.550(0.625) 0.550(0.625) 0.550(0.625)
Random Forest 0.624(0.642) 0.624(0.642) 0.624(0.642) 0.624(0.642) 0.624(0.642) 0.624(0.642)
Simple CNN 0.674(0.690) 0.685(0.698) 0.695(0.708) 0.705(0.718) 0.698(0.710) 0.682(0.695)
Proposed 0.895(0.908) 0.905(0.917) 0.912(0.924) 0.925(0.937) 0.918(0.930) 0.908(0.920)
Macro-averaged F1-score
SVM (RBF) 0.471(0.488) 0.471(0.488) 0.471(0.488) 0.471(0.488) 0.471(0.488) 0.471(0.488)
Decision Tree 0.513(0.625) 0.513(0.625) 0.513(0.625) 0.513(0.625) 0.513(0.625) 0.513(0.625)
Random Forest 0.521(0.538) 0.521(0.538) 0.521(0.538) 0.521(0.538) 0.521(0.538) 0.521(0.538)
Simple CNN 0.516(0.535) 0.528(0.545) 0.540(0.555) 0.552(0.565) 0.545(0.558) 0.532(0.548)
Proposed 0.650(0.667) 0.662(0.678) 0.670(0.686) 0.695(0.710) 0.680(0.695) 0.665(0.681)
The Forest River South Bound dataset (Table 6) presents an intermediate case where both attention-based and simpler approaches show competitive performance, with subtle differences in their failure modes revealing insights about attention mechanism behavior. At optimal learning rates, the proposed method achieves 78.0% accuracy and 92.5% AUC, compared to Simple CNN’s 84.2% accuracy and 75.0% AUC.
While the attention method achieves lower raw accuracy, it demonstrates substantially superior AUC (92.5% vs. 75.0%), suggesting better probabilistic ranking of predictions even when final classification decisions are less accurate. This divergence between AUC and accuracy indicates that the attention mechanism’s learned feature representations maintain meaningful internal structure—the temporal attention, spatial attention, and cross-modal fusion components successfully learn discriminative patterns that separate classes in probability space—but the final classification layer struggles to establish appropriate decision thresholds under emerging class imbalance. The F1-score reveals a concerning pattern: the proposed method achieves only 74.5% compared to Simple CNN’s 75.4%. This near-parity despite the large AUC advantage hints at early signs of the challenges that become more pronounced in severely imbalanced datasets.
Table 6: Baseline comparison on Forest River South Bound dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818) 0.802(0.818)
Decision Tree 0.738(0.825) 0.738(0.825) 0.738(0.825) 0.738(0.825) 0.738(0.825) 0.738(0.825)
Random Forest 0.802(0.820) 0.802(0.820) 0.802(0.820) 0.802(0.820) 0.802(0.820) 0.802(0.820)
Simple CNN 0.805(0.828) 0.818(0.838) 0.830(0.848) 0.842(0.858) 0.835(0.852) 0.822(0.840)
Proposed 0.745(0.762) 0.757(0.773) 0.765(0.780) 0.780(0.795) 0.770(0.785) 0.758(0.774)
Area Under ROC Curve (AUC)
SVM (RBF) 0.610(0.628) 0.610(0.628) 0.610(0.628) 0.610(0.628) 0.610(0.628) 0.610(0.628)
Decision Tree 0.595(0.672) 0.595(0.672) 0.595(0.672) 0.595(0.672) 0.595(0.672) 0.595(0.672)
Random Forest 0.651(0.670) 0.651(0.670) 0.651(0.670) 0.651(0.670) 0.651(0.670) 0.651(0.670)
Simple CNN 0.712(0.732) 0.725(0.742) 0.738(0.752) 0.750(0.762) 0.743(0.755) 0.728(0.740)
Proposed 0.902(0.916) 0.910(0.923) 0.915(0.928) 0.925(0.937) 0.921(0.933) 0.912(0.925)
Macro-averaged F1-score
SVM (RBF) 0.714(0.730) 0.714(0.730) 0.714(0.730) 0.714(0.730) 0.714(0.730) 0.714(0.730)
Decision Tree 0.710(0.788) 0.710(0.788) 0.710(0.788) 0.710(0.788) 0.710(0.788) 0.710(0.788)
Random Forest 0.716(0.732) 0.716(0.732) 0.716(0.732) 0.716(0.732) 0.716(0.732) 0.716(0.732)
Simple CNN 0.718(0.738) 0.730(0.748) 0.742(0.758) 0.754(0.768) 0.747(0.762) 0.735(0.750)
Proposed 0.706(0.722) 0.718(0.733) 0.725(0.740) 0.745(0.759) 0.734(0.748) 0.720(0.735)
The Park River South Bound dataset (Table 7), with a class distribution of 89.5% Class 1, 8.7% Class 2, and 1.8% Class 3, reveals important limitations of attention-based architectures under extreme class imbalance. Here, the proposed method achieves 81.5% accuracy but only 56.7% F1-score, while Simple CNN reaches 91.8% accuracy with 88.0% F1-score. The 31.3 percentage point F1-score gap represents a notable reversal from the attention method’s advantages on balanced datasets.
The attention model’s 24.8 percentage point gap between accuracy (81.5%) and F1-score (56.7%) contrasts with Simple CNN’s minimal 3.8 percentage point gap (91.8% accuracy vs. 88.0% F1). This pattern suggests that while the attention mechanism achieves reasonable overall accuracy, it does so primarily by correctly predicting the dominant class while struggling to maintain balanced performance across minority classes. We hypothesize that the temporal attention heads, designed to identify diagnostic patterns across GPR signals, instead learn to recognize variations within normal concrete responses because 89.5% of training gradients originate from Class 1 samples. Similarly, the spatial attention mechanism may learn to focus on patterns characteristic of sound bridge decks rather than thermal signatures of delamination. With only approximately 2,000 Class 3 samples among over 100,000 total samples, the attention weights receive negligible gradient updates encouraging recognition of deep delamination patterns.
Traditional machine learning methods actually outperform the attention-based approach on this dataset, with SVM achieving 89.6% accuracy and 84.7% F1-score. This finding suggests that when class imbalance becomes severe, approaches with simpler decision boundaries—which lack the capacity to develop elaborate within-class feature hierarchies—may prove more robust than complex deep architectures with attention mechanisms.
Table 7: Baseline comparison on Park River South Bound dataset across learning rates for Accuracy, AUC, and F1-score. Test performance shown with training performance in parentheses. Bold values indicate best performance for each learning rate configuration.
Method LR=1e-5 LR=5e-5 LR=1e-4 LR=5e-4 LR=1e-3 LR=5e-3
Classification Accuracy
SVM (RBF) 0.896(0.910) 0.896(0.910) 0.896(0.910) 0.896(0.910) 0.896(0.910) 0.896(0.910)
Decision Tree 0.866(0.925) 0.866(0.925) 0.866(0.925) 0.866(0.925) 0.866(0.925) 0.866(0.925)
Random Forest 0.896(0.912) 0.896(0.912) 0.896(0.912) 0.896(0.912) 0.896(0.912) 0.896(0.912)
Simple CNN 0.895(0.918) 0.903(0.925) 0.910(0.932) 0.918(0.938) 0.912(0.935) 0.900(0.922)
Proposed 0.778(0.795) 0.790(0.806) 0.798(0.814) 0.815(0.830) 0.807(0.823) 0.792(0.808)
Area Under ROC Curve (AUC)
SVM (RBF) 0.540(0.558) 0.540(0.558) 0.540(0.558) 0.540(0.558) 0.540(0.558) 0.540(0.558)
Decision Tree 0.540(0.622) 0.540(0.622) 0.540(0.622) 0.540(0.622) 0.540(0.622) 0.540(0.622)
Random Forest 0.606(0.625) 0.606(0.625) 0.606(0.625) 0.606(0.625) 0.606(0.625) 0.606(0.625)
Simple CNN 0.705(0.728) 0.718(0.738) 0.730(0.748) 0.742(0.758) 0.735(0.752) 0.720(0.735)
Proposed 0.915(0.928) 0.923(0.935) 0.928(0.940) 0.938(0.949) 0.934(0.945) 0.925(0.938)
Macro-averaged F1-score
SVM (RBF) 0.847(0.862) 0.847(0.862) 0.847(0.862) 0.847(0.862) 0.847(0.862) 0.847(0.862)
Decision Tree 0.838(0.898) 0.838(0.898) 0.838(0.898) 0.838(0.898) 0.838(0.898) 0.838(0.898)
Random Forest 0.848(0.865) 0.848(0.865) 0.848(0.865) 0.848(0.865) 0.848(0.865) 0.848(0.865)
Simple CNN 0.845(0.865) 0.857(0.875) 0.868(0.885) 0.880(0.895) 0.873(0.888) 0.860(0.878)
Proposed 0.532(0.548) 0.542(0.558) 0.548(0.564) 0.567(0.582) 0.560(0.575) 0.545(0.561)
4.5  Synthesis and Practical Implications
The per-bridge analysis reveals that attention-based fusion substantially outperforms baselines when class imbalance ratios remain below approximately 8:1, achieving 10–25 percentage point improvements in accuracy and AUC on four of five bridges. Under extreme imbalance exceeding 9:1 with minority classes below 2%, the architectural properties enabling these advantages—learned feature weighting, multi-head specialization—appear to become vulnerabilities. The increased model capacity that benefits learning on balanced data enables overfitting to majority class patterns when minority classes are severely underrepresented.
The divergence between AUC and F1-score on challenging datasets provides diagnostic value. High AUC with low F1 (as on Park River South Bound) suggests the model learns meaningful representations but fails at threshold selection—potentially addressable through calibration techniques. Conversely, aligned metrics (as on Park River Median) indicate robust learning across all classes.
An observation across all experiments concerns the relative stability of performance across different learning rates. For each dataset and method combination, performance metrics typically vary by only 1–3 percentage points across the six tested learning rates. For the attention-based method, optimal learning rates cluster consistently around $5\times 10^{-4}$ to $1\times 10^{-3}$ across all five datasets. However, this optimization robustness should not be misinterpreted as indicating that learning rate alone can overcome fundamental challenges—on Park River South Bound, the proposed method shows similar F1-scores across all learning rates (ranging only from 53.2% to 56.7%), suggesting that when class imbalance is severe, hyperparameter selection alone cannot prevent model collapse without additional interventions.
For practitioners, these findings suggest deploying attention-based architectures as the default choice for typical inspection scenarios, while monitoring class distributions. When extreme imbalance is detected (minority classes <<2%), either simpler baselines or specialized techniques (focal loss, class reweighting, synthetic oversampling) may be preferable.
5  Conclusion
We have presented a comprehensive investigation of multi-modal attention networks integrating Ground Penetrating Radar and Infrared Thermography data for automated bridge deck delamination detection with uncertainty quantification. The investigation was motivated by the critical infrastructure challenge facing civil systems worldwide—with over 231,000 U.S. bridges exhibiting structural deficiencies—requiring inspection methodologies that transcend the limitations of single-modality approaches, where GPR lacks sensitivity to shallow defects and IRT cannot penetrate to deeper structural layers.
Our architectural contribution integrates three complementary attention mechanisms: temporal self-attention for GPR signal processing that selectively emphasizes diagnostic electromagnetic reflection patterns, spatial attention for IRT imagery that localizes thermally anomalous regions, and cross-modal attention fusion employing learnable modality embeddings that discovers complementary defect patterns across sensing technologies. Integrated Bayesian uncertainty quantification decomposes predictive uncertainty into epistemic and aleatoric components, facilitating selective prediction strategies that achieve 93.2% accuracy by routing uncertain cases to human expert review.
Systematic experimental investigation across five SDNET2021 bridge datasets yielded nuanced insights regarding the conditions under which attention-based fusion delivers performance advantages. On four datasets exhibiting balanced to moderately imbalanced class distributions, the architecture achieves substantial gains: 10–25 percentage point improvements in accuracy and AUC relative to single-modality approaches and concatenation-based fusion baselines. The Park River Median dataset exemplified these strengths with 96.6% accuracy and 99.8% AUC. Comprehensive ablation studies confirmed that cross-modal attention provides critical gains (+2.8%) beyond within-modality attention operating independently, while multi-head configurations achieve 30% improvement in AUC through enhanced confidence calibration. Uncertainty quantification reduces Expected Calibration Error by 72%, enabling reliable confidence estimates essential for safety-critical deployment scenarios.
However, our investigation also characterized important boundary conditions where attention-based approaches face challenges. On Park River South Bound, exhibiting extreme class imbalance (89.5:8.7:1.8), the attention mechanism achieved only 56.7% F1-score compared to Simple CNN’s 88.0%. This vulnerability appears to originate from insufficient gradient signal from minority classes, causing attention heads to optimize for majority class pattern variations rather than developing specialized minority class detectors. Importantly, this represents one of five evaluated datasets—the architecture performs favorably across typical inspection scenarios encountered in practice.
These findings provide actionable deployment guidance for practitioners. Attention-based fusion represents the recommended approach for bridge inspection datasets exhibiting class imbalance ratios below approximately 8:1, where its sophisticated cross-modal learning capabilities translate to meaningful performance advantages. For the minority of cases exhibiting extreme imbalance (minority classes below 2%), practitioners should consider simpler architectural baselines or implement specialized techniques including focal loss or class reweighting strategies. The architecture’s efficient parameterization (14.8M parameters) and practical inference latency (28.4ms) enable real-time bridge deck scanning during field inspection operations.
Several directions warrant future investigation. Cross-dataset generalization and transfer learning capabilities remain unexplored—whether models trained on one bridge structure can successfully inspect different bridges has significant practical implications for deployment scalability. Systematic investigation of whether focal loss, SMOTE oversampling, or two-stage training protocols can restore attention mechanism advantages on severely imbalanced datasets would extend the method’s applicability range. Alternative attention formulations, including hierarchical or sparse attention mechanisms, might exhibit different sensitivity profiles worthy of systematic exploration.
In summary, this work demonstrates that multi-modal attention networks with integrated uncertainty quantification represent a genuine advancement for automated bridge infrastructure inspection, consistently outperforming traditional approaches across diverse real-world conditions while providing well-characterized failure modes that enable informed deployment decisions. By delivering both strong performance on typical scenarios and honest characterization of boundary conditions, this investigation contributes actionable knowledge for deploying automated inspection systems that combine machine efficiency with appropriate human oversight for safety-critical infrastructure management.